### v1

# Exact penalty functions in optimization with unbounded constraint sets

Abstract.
This paper identifies necessary and sufficient conditions for the exactness of penalty functions in optimization problems whose constraint sets are not necessarily bounded. The case where the data of problems is locally Lipschitz or semi-algebraic is studied in detail. The conditions are given in terms of properties of the objective and residual functions of the problems in question. The obtained results generalize and improve some known results in the literature on exact penalty functions.
Key words and phrases:
constrained optimization, exact penalty, calmness, semi-algebraic, Łojasiewicz inequality
2010 Mathematics Subject Classification:
Primary 90C26; Secondary 49J52, 14P10
The work was partially supported by the Chinese National Natural Science Foundation under grant numbers 12471478, 12371300.
Corresponding author
1. Introduction
Constrained optimization problems arise in various fields, including engineering, economics, and machine learning, where the goal is to optimize an objective function subject to a set of constraints. Problems with complicated constraints are very difficult to deal with. A common approach for solving such problems is the usage of penalty function methods, wherein a constrained optimization problem is transformed into a sequence of unconstrained optimization problems whose solutions ideally converge to a solution of the original constrained problem; the unconstrained problems are formed by bringing the constraints into the objective function via a residual function and a scalar penalty parameter. Among these methods, exact penalty functions are particularly important because they allow us to solve a constrained optimization problem by solving a single unconstrained optimization problem. More precisely, a penalty function is said to be exact (or to have the exact penalty property) if there is a penalty parameter for which the corresponding unconstrained optimization problem and the original constrained optimization problem have the same optimal value and the same optimal solutions.
The concept of exact penalization traces back to the seminal works of Eremin [15] and Zangwill [36] (see also [30]). There is a huge literature on all aspects of the theory and applications of exact penalty functions; for more details, we refer the reader to the monographs [9, 16, 20, 21, 27, 29, 33, 38], the surveys [4, 6, 17] and the recent papers [18, 22, 31, 35] with the references therein.
In this paper we would like to identify necessary and sufficient conditions for a penalty function to be exact; the conditions are given in terms of properties of the objective and residual functions of the problems in question. As far as we know, for optimization problems whose constraint sets are unbounded, there are not many such studies; see [2, 8, 13, 37, 38, 39].
Contributions
We mainly study the exactness of penalty functions for constrained optimization problems with potentially unbounded constraint sets. Among other things, our main contributions are given in several steps:
  • A necessary and sufficient condition for a penalty function to be exact is given (see Theorem 3.1).
  • An equivalent relation between the exactness of a penalty function and the global calmness for the problem in question is established (see Theorem 3.2); this relation can be seen as a global version of [7, Proposition 6.4.3] and [5, Theorem 1.1].
  • For problems defined by locally Lipschitz functions, under some regularity conditions, a new necessary and sufficient condition for a penalty function to be exact is derived (see Theorem 4.1); this result improves [8, Theorem 4.2] and [37, Theorems 1.5 and 3.1].
  • Based on the study of Łojasiewicz inequalities on unbounded sets, exact penalty functions for constrained optimization problems with semi-algebraic data are proposed. In the semi-algebraic setting, these results generalize [34, Theorem 1], [10, Theorem 3.1] and [27, Theorem 2.1.2].
At this point, we would like to note the following facts:
  • the closeness of constraint sets and the continuity of objective functions, except for Theorem 4.1, are not required;
  • the boundedness of constraint sets is not imposed; and
  • the existence of optimal solutions is not assumed.
We confine our study to the finite-dimensional case for two reasons. First of all, to lighten the exposition, we would like to concentrate on the basic ideas without technical and notational complications. Furthermore, the results obtained in Section 5 are peculiar to finite-dimensions.
The rest of the paper is organized as follows. Some definitions and preliminary results from variational analysis and semi-algebraic geometry are recalled in Section 2. The main results and their proofs are presented in Sections 3, 4 and 5. Conclusions are given in Section 6.
2. Preliminaries
2.1. Notation
[…]
Let $\mathbb{R}_{+}:=[0,\infty),\mathbb{R}_{-}:=(-\infty,0]$ and $\overline{\mathbb{R}}:=\mathbb{R}\cup\{\infty\}.$ For a real number $r,$ we write $[r]_{+}:=\max\{r,0\}.$
For a nonempty set $\Omega\subset\mathbb{R}^{n},$ the closure, convex hull and conic hull of $\Omega$ are denoted, respectively, by $\mathrm{cl}\,{\Omega},$ $\mathrm{co}\,\Omega$ and $\mathrm{cone}\,\Omega.$ We will associate with $\Omega$ the distance function
$\mathrm{dist}(\cdot,\Omega)\colon\mathbb{R}^{n}\to\mathbb{R},\quad x\mapsto\inf_{x^{\prime}\in\Omega}\|x-x^{\prime}\|,$
and define the Euclidean projector of $x\in\mathbb{R}^{n}$ to $\Omega$ by
$\Pi_{\Omega}(x)$ ::=:= $\{x^{\prime}\in\Omega\ |\ \|x-x^{\prime}\|=\mathrm{dist}(x,\Omega)\}.$
The indicator function $\delta_{\Omega}\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$ of $\Omega$ is defined by
$\delta_{\Omega}(x)$ ::=:= $\begin{cases}0&\textrm{ if }x\in\Omega,\\ \infty&\textrm{ otherwise.}\end{cases}$
For an extended real-valued function $f\colon\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}},$ we denote its effective domain, graph, and epigraph by, respectively,
$\mathrm{dom}f$ ::=:= $\{x\in\mathbb{R}^{n}\ |\ f(x)<\infty\},$
$\mathrm{gph}f$ ::=:= $\{(x,y)\in\mathbb{R}^{n}\times\mathbb{R}\ |\ f(x)=y\},$
$\mathrm{epi}f$ ::=:= $\{(x,y)\in\mathbb{R}^{n}\times\mathbb{R}\ |\ f(x)\leq y\}.$
We call $f$ a proper function if $f(x)<\infty$ for at least one $x\in\mathbb{R}^{n},$ or in other words, if $\mathrm{dom}f$ is a nonempty set. The function $f$ is said to be lower semi-continuous if its epigraph is a closed set.
For an extended real-valued function $f\colon\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ and a set $\Omega\subset\mathbb{R}^{n},$ we let
$\inf_{\Omega}f$ ::=:= $\inf_{x\in\Omega}f(x)\ :=\ \inf\{f(x)\mid x\in\Omega\}.$
We introduce notation also for the sets of points $x$ where the minimum of $f$ over $\Omega$ is regarded as being attained:
$\mathrm{argmin}_{\Omega}f(x)$ ::=:= $\begin{cases}\{x\in\Omega\mid f(x)=\inf_{x\in\Omega}f(x)\}&\textrm{ if }\inf_{x\in\Omega}f(x)\neq\infty,\\ \emptyset&\textrm{ otherwise.}\end{cases}$
For a set-valued mapping $F\colon\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m},$ we denote its graph by
$\mathrm{gph}F$ ::=:= $\{(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{m}\mid y\in F(x)\};$
the Painlevé–Kuratowski outer limit of $F$ at $x\in\mathbb{R}^{n}$ is defined by
$\mathop{{\rm Lim}\,{\rm sup}}_{x^{\prime}\to{x}}F(x)$ ::=:= $\{y\in\mathbb{R}^{m}\mid\exists x_{k}\to{x},\exists y_{k}\in F(x_{k}),y_{k}\to y\}.$
2.2. Normal cones, subdifferentials and coderivatives
Here we recall some definitions and results of variational analysis, which can be found in [28, 29, 33].
Definition 2.1.
Consider a set $\Omega\subset\mathbb{R}^{n}$ and a point ${x}\in\Omega.$ The limiting normal cone (also known as the basic or Mordukhovich normal cone) to $\Omega$ at ${x}$ is defined by
$N_{\Omega}(x)$ ::=:= $\mathop{{\rm Lim}\,{\rm sup}}_{x^{\prime}\rightarrow x}\Big{[}\mathrm{cone}\big{(}x^{\prime}-\Pi_{\Omega}(x^{\prime})\big{)}\Big{]}.$
If $x\not\in\Omega,$ we put ${N}_{\Omega}({x}):=\emptyset.$
Remark 2.1.
(i) It is well-known that $N_{\Omega}({x})$ is a closed (possibly non-convex) cone.
(ii) If $\Omega$ is a manifold of class $C^{1},$ then the normal cone $N_{\Omega}({x})$ is equal to the normal space to $\Omega$ at ${x}$ in the sense of differential geometry; see [33, Example 6.8].
Definition 2.2.
Consider an extended real-valued function $f\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$ and a point ${x}\in\mathrm{dom}f.$ The limiting (or Mordukhovich) subdifferential of $f$ at ${x}$ is defined by
$\partial f({x})$ ::=:= $\{u\in\mathbb{R}^{n}\mid(u,-1)\in{N}_{\mathrm{epi}f}({x},f({x}))\}.$
If $x\not\in\mathrm{dom}f,$ we put $\partial f({x}):=\emptyset.$
Remark 2.2.
In [28, 29, 33] the reader can find equivalent analytic descriptions of the limiting subdifferential $\partial f({x})$ and comprehensive studies of it and related constructions. For convex $f,$ this subdifferential coincides with the convex subdifferential. Furthermore, if the function $f$ is of class $C^{1},$ then $\partial f({x})=\{\nabla f({x})\}.$
Definition 2.3.
Consider a set-valued mapping $F\colon\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ and a point $({x},{y})\in\mathrm{gph}F.$ The (basic) coderivative of $F$ at $({x},{y})$ is the set-valued mapping $D^{*}F({x},{y})\colon\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{n}$ defined by
$D^{*}F({x},{y})(v)$ ::=:= $\{u\in\mathbb{R}^{n}\mid(u,-v)\in N_{\mathrm{gph}F}({x},{y})\}\quad\textrm{ for all }\quad v\in\mathbb{R}^{m}.$
The following facts are well-known (see, for example, [29]).
Lemma 2.1.
Consider a nonempty closed set $\Omega\subset\mathbb{R}^{n}.$ We have for all ${x}\in\Omega,$
$N_{\Omega}({x})$ $=$ $\mathop{{\rm Lim}\,{\rm sup}}_{x^{\prime}\xrightarrow{\Omega}{x}}{N}_{\Omega}(x^{\prime}),$
where $x^{\prime}\xrightarrow{\Omega}{x}$ means that $x^{\prime}\rightarrow{x}$ with $x^{\prime}\in\Omega.$
Lemma 2.2.
Let $\Omega\subset\mathbb{R}^{n}$ be a nonempty closed set. Then
$\partial\mathrm{dist}(\cdot,\Omega)(x)$ $=$ $\begin{cases}N_{\Omega}(x)\cap\mathbb{B}&\quad\textrm{ if }x\in\Omega,\\ \frac{x-\Pi_{\Omega}(x)}{\mathrm{dist}(x,\Omega)}&\quad\textrm{ otherwise.}\end{cases}$
Lemma 2.3.
For a lower semicontinuous function $f\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$ and a point $\overline{x}\in\mathrm{dom}f,$ we have
[…]
Lemma 2.4.
Let $f_{i}\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$, $i=1,\dots,m$ with $m\geq 2$, be lower semi-continuous at $\overline{x}\in\mathbb{R}^{n}$ and let all but one of these functions be Lipschitz continuous around $\overline{x}.$ Then the following inclusions hold
$\partial\left(f_{1}+\cdots+f_{m}\right)(\overline{x})$$\subset$ $\partial f_{1}(\overline{x})+\cdots+\partial f_{m}(\overline{x}),$
$\partial(\max f_{i})(\overline{x})$$\subset$ $\bigcup\mathrm{co}(\{v_{i}\mid i\in I(\overline{x})\}),$
where $I(\overline{x}):=\{i\in\{1,\ldots,m\}\mid f_{i}(\overline{x})=\max_{j}f_{j}(\overline{x})\}$ and the union is taken over all vectors $v_{i}\in\partial f_{i}(\overline{x})$ for $i\in I(\overline{x}).$
Lemma 2.5 (Fermat rule).
If a proper function $f\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$ has a local minimum at $\overline{x}\in\mathrm{dom}f,$ then $0\in\partial f(\overline{x}).$
Lemma 2.6.
Let $f\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be Lipschitz continuous around $\overline{x}\in\mathbb{R}^{n}$ with modulus $L.$ Then $\partial f(\overline{x})\neq\emptyset$ and
$\|v\|$$\leq$ $L\quad\textrm{ for all }\quad v\in\partial f(\overline{x}).$
Lemma 2.7 (chain rule).
Let $g\colon\mathbb{R}^{n}\to{\mathbb{R}}^{m}$ be Lipschitz continuous around $\overline{x}\in\mathbb{R}^{n}$ and $h\colon\mathbb{R}^{m}\to{\mathbb{R}}$ be Lipschitz continuous around $\overline{y}:=g(\overline{x})\in\mathbb{R}^{m}.$ Then
$\partial(h\circ g)(\overline{x})$$\subset$ $\bigcup_{w\in\partial h(\overline{y})}\partial\langle w,g\rangle(\overline{x}).$
Finally, the next lemma expresses the coderivative of a single-valued Lipschitz mapping $g\colon\mathbb{R}^{n}\to{\mathbb{R}}^{m}$ via the subdifferential of the scalarization $\langle w,g\rangle\colon\mathbb{R}^{n}\to{\mathbb{R}},x\mapsto\langle w,g(x)\rangle.$
Lemma 2.8.
Let $g\colon\mathbb{R}^{n}\to{\mathbb{R}}^{m}$ be Lipschitz continuous around $\overline{x}\in\mathbb{R}^{n}.$ Then
$D^{*}g(\overline{x})(w)$ $=$ $\partial\langle w,g\rangle(\overline{x})\quad\textrm{ for all }\quad w\in\mathbb{R}^{m}.$
2.3. Semi-algebraic geometry
Now, we recall some notions and results of semi-algebraic geometry, which can be found in [1, 3, 19].
Definition 2.4.
A subset $S$ of $\mathbb{R}^{n}$ is semi-algebraic if it is a finite union of sets of the form
$\{x\in\mathbb{R}^{n}\ |\ f_{i}(x)=0,\ i=1,\ldots,k;f_{j}(x)>0,\ j=k+1,\ldots,p\},$
where all $f_{i}$ are polynomials. In other words, $S$ is a union of finitely many sets, each defined by finitely many polynomial equalities and inequalities.
A function $f\colon S\rightarrow{\mathbb{R}\cup\{\infty\}}$ is said to be semi-algebraic if its graph
$\{(x,y)\in S\times\mathbb{R}\ |\ y=f(x)\}$
is a semi-algebraic set.
Example 2.1.
Each polyhedral set is the intersection of a finite number of half-spaces, and so is semi-algebraic.
A major fact concerning the class of semi-algebraic sets is its stability under linear projections.
Theorem 2.1 (Tarski–Seidenberg theorem).
The image of any semi-algebraic set $S\subset\mathbb{R}^{n}$ under a projection to any linear subspace of $\mathbb{R}^{n}$ is a semi-algebraic set.
Remark 2.3.
As an immediate consequence of the Tarski–Seidenberg theorem, we get the semi-algebraicity of any set $\{x\in A\ |\ \exists y\in B,(x,y)\in C\},$ provided that $A,B,$ and $C$ are semi-algebraic sets in the corresponding spaces. Also, $\{x\in A\ |\ \forall y\in B,(x,y)\in C\}$ is a semi-algebraic set as its complement is the union of the complement of $A$ and the set $\{x\in A\ |\ \exists y\in B,(x,y)\not\in C\}.$ Thus, if we have a finite collection of semi-algebraic sets, then any set obtained from them with the help of a finite chain of quantifiers on variables is also semi-algebraic.
The following well known lemmas will be of great importance for us (see, for example, [1, 3, 19]).
Lemma 2.9.
Every semi-algebraic set in $\mathbb{R}$ is a finite union of points and intervals.
Lemma 2.10 (growth dichotomy lemma).
  1. (i)
    Let $f\colon(0,\epsilon)\rightarrow{\mathbb{R}}$ be a semi-algebraic function with $f(t)\neq 0$ for all $t\in(0,\epsilon).$ Then there exist constants $a\neq 0$ and $\alpha\in{\mathbb{Q}}$ such that $f(t)=at^{\alpha}+o(t^{\alpha})$ as $t\to 0^{+}.$
  2. (ii)
    Let $f\colon(r,+\infty)\rightarrow{\mathbb{R}}$ be a semi-algebraic function with $f(t)\neq 0$ for all $t\in(r,+\infty).$ Then there exist constants $b\neq 0$ and $\beta\in{\mathbb{Q}}$ such that $f(t)=bt^{\beta}+o(t^{\beta})$ as $t\to+\infty.$
Lemma 2.11 (monotonicity lemma).
Let $f\colon(a,b)\to\mathbb{R}$ be a semi-algebraic function. Then there are $a=a_{0}<a_{1}<\cdots<a_{s}<a_{s+1}=b$ such that, for each $i=0,\ldots,s,$ the restriction $f|_{(a_{i},a_{i+1})}$ is analytic and either constant, strictly increasing, or strictly decreasing.
3. Exact penalty functions: the general case
Throughout this paper, we consider the following constrained optimization problem
minimize $\textrm{minimize }f(x)\quad\textrm{ subject to }x\in S,$ (P)
where $S$ is a nonempty subset of $\mathbb{R}^{n}$ and $f\colon\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is an extended real-valued function. To avoid triviality, we assume in what follows that the optimal value $f_{*}:=\inf_{x\in S}f(x)$ of the problem (P) is finite. In this section, we will present some global necessary and sufficient condition for a penalty function to be exact. Let us start with the following.
Definition 3.1.
A residual function of $S$ is a nonnegative valued function $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ such that $\psi(x)=0$ if and only if $x\in S.$
Remark 3.1.
If $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ is a residual function for $S,$ then for any function $\phi\colon\mathbb{R}^{n}\to(0,\infty)$ and any constants $\alpha>0$ and $\beta>0,$ one has the functions $\phi\cdot\psi,$ $\psi^{\alpha}$ and $\psi^{\alpha}+\psi^{\beta}$ are residual for $S.$
The following simple fact will play an important role in establishing exact penalty properties in optimization problems with unbounded constraint sets.
Theorem 3.1.
For any residual function $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ of $S,$ the following three properties are equivalent:
  1. (i)
    for all $c$ sufficiently large, one has
    $\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c\,\psi(x)\big{\}};$
  2. (ii)
    there exists a constant $c_{*}>0$ such that
    $\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c_{*}\,\psi(x)\big{\}};$
  3. (iii)
    there exists a constant $c_{*}>0$ such that
    $c_{*}\psi(x)$$\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad x\in\mathbb{R}^{n}.$ (1)
When these equivalent properties hold, one has for all $c>c_{*}$ that
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big{\{}f(x)+c\,\psi(x)\big{\}}.$
Proof.
The implications (i) $\Rightarrow$ (ii) $\Rightarrow$ (iii) is obvious.
(iii) $\Rightarrow$ (i) Fix $c>c_{*}$ and take any $x\in\mathbb{R}^{n}.$ If $f(x)\geq f_{*},$ then $f(x)+c\,\psi(x)\geq f_{*};$ otherwise (1) gives
$f(x)+c\,\psi(x)$$\geq$ $f(x)+c_{*}\,\psi(x)\ \geq\ f(x)+[f_{*}-f(x)]_{+}\ =\ f_{*}.$
In both cases we have
$f(x)+c\,\psi(x)$$\geq$ $f_{*}.$
Since $x\in\mathbb{R}^{n}$ was chosen arbitrarily, we conclude that
$\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c\,\psi(x)\big{\}}$$\geq$ $f_{*}.$
On the other hand, by definition, we have
$\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c\,\psi(x)\big{\}}$$\leq$ $\inf_{x\in S}\big{\{}f(x)+c\,\psi(x)\big{\}}\ =\ \inf_{x\in S}f(x)\ =\ f_{*}.$
Therefore, (i) holds.
To show the last conclusion, assume that $x_{0}$ minimizes $f$ on $S.$ Then
$f(x_{0})+c\,\psi(x_{0})$ $=$ $f(x_{0})\ =\ f_{*},$
which, together with (i), yields that $x_{0}$ minimizes $f+c\psi$ on $\mathbb{R}^{n}.$
Conversely, assume that $x_{0}$ minimizes $f+c\psi$ on $\mathbb{R}^{n}.$ Since $c>c_{*}$ and the function $\psi$ is nonnegative, it holds that
$f(x_{0})+c_{*}\psi(x_{0})\leq f(x_{0})+c\,\psi(x_{0})$$\leq$ $f_{*}.$
On the other hand, it follows from (1) that
$f(x_{0})+c_{*}\psi(x_{0})$$\geq$ $f_{*}.$
Therefore,
$f(x_{0})+c_{*}\psi(x_{0})$ $=$ $f(x_{0})+c\,\psi(x_{0})\ =\ f_{*}.$
Since $c_{*}<c,$ then $\psi(x_{0})=0.$ Hence $x_{0}\in S$ and $f(x_{0})=f_{*}.$ Thus, $x_{0}$ minimizes $f$ on $S.$
Example 3.1.
Let $n:=1,S:=\{x\in\mathbb{R}\mid x<1\}$ and $f(x):=|x|+\delta_{\{x\neq 0\}}(x).$ By definition, the function $\psi\colon\mathbb{R}\to\mathbb{R}$ defined by
$\psi(x):=\begin{cases}\max\{x-1,0\}&\textrm{ if }x\neq 1,\\ 1&\textrm{ otherwise}\end{cases}$
is a residual function for $S.$ Fix any $c_{*}>0.$ It is easy to check that $f_{*}=0$ and
$c_{*}\psi(x)$$\geq$ $0\ =\ [f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad x\in\mathbb{R}.$
By Theorem 3.1, for any $c>c_{*}$ we have
$\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}}\big{\{}f(x)+c\,\psi(x)\big{\}},$
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}}\big{\{}f(x)+c\,\psi(x)\big{\}}\ =\ \emptyset.$
Also observe that the constraint set $S$ is neither closed nor bounded and the objective function $f$ is not lower semicontinuous.
The results presented in the rest of this section are inspired by the works of Clarke [7, Section 6.4] and Burke [5, 6]. To see this, assume that
$S$ ::=:= $\{x\in\mathbb{R}^{n}\mid g(x)\in C\},$
where $g\colon\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is a mapping and $C$ is a nonempty subset of $\mathbb{R}^{m}.$ Then the problem (P) can be rewritten as
minimize $\textrm{minimize }f(x)\quad\textrm{ subject to }g(x)\in C.$ (P0)
In this case, an equivalent relation between the exact penalization and the global calmness can be stated as follows.
Theorem 3.2.
The following two properties are equivalent:
  1. (i)
    for all $c$ sufficiently large one has
    $\inf_{g(x)\in{C}}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c\,\mathrm{dist}\big{(}g(x),{C}\big{)}\big{\}};$
  2. (ii)
    the problem (P0) is globally calm in the sense that there exists a constant $c_{*}>0$ such that for every pair $(\overline{x},\overline{u})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ with $g(\overline{x})\in{C}+\overline{u},$ one has
    $\inf_{g(x)\in{C}}f(x)$$\leq$ $f(\overline{x})+c_{*}\,\|\overline{u}\|.$
When these equivalent properties hold, one has for all $c>c_{*}$ that
$\mathrm{argmin}_{g(x)\in C}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big{\{}f(x)+c\,\mathrm{dist}(g(x),C)\big{\}}.$
Proof.
It suffices to show the equivalence of (i) and (ii) since the last conclusion is a direct consequence of Theorem 3.1.
(i) $\Rightarrow$ (ii) Assume that
$\inf_{g(x)\in{C}}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c_{*}\,\mathrm{dist}\big{(}g(x),{C}\big{)}\big{\}}$
for some $c_{*}>0.$ Let $(\overline{x},\overline{u})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ be such that $g(\overline{x})\in{C}+\overline{u}.$ Then
$\inf_{g(x)\in{C}}f(x)$$\leq$ $f(\overline{x})+c_{*}\,\mathrm{dist}\big{(}g(\overline{x}),{C}\big{)}$
$=$ $f(\overline{x})+c_{*}\,\inf\big{\{}\|g(\overline{x})-y\|\mid y\in{C}\big{\}}$
$\leq$ $f(\overline{x})+c_{*}\,\inf\big{\{}\|g(\overline{x})-(y+{\overline{u}})\|+\|\overline{u}\|\mid y\in{C}\big{\}}$
$=$ $f(\overline{x})+c_{*}\,\mathrm{dist}\big{(}g(\overline{x}),{C}+{\overline{u}}\big{)}+c_{*}\,\|\overline{u}\|$
$=$ $f(\overline{x})+c_{*}\,\|\overline{u}\|,$
which gives the desired result.
(ii) $\Rightarrow$ (i) Let $\overline{x}\in\mathbb{R}^{n}$ and take any $\epsilon>0.$ There exists a point $y\in{C}$ such that
$\|g(\overline{x})-y\|$$\leq$ $\mathrm{dist}\big{(}g(\overline{x}),{C}\big{)}+\epsilon.$
Let $\overline{u}:=g(\overline{x})-y.$ Then $g(\overline{x})=y+\overline{u}\in{C}+\overline{u}.$ By hypothesis, we get
$\inf_{g(x)\in{C}}f(x)$$\leq$ $f(\overline{x})+c_{*}\,\|\overline{u}\|$
$\leq$ $f(\overline{x})+c_{*}\,\mathrm{dist}\big{(}g(\overline{x}),{C}\big{)}+c_{*}\epsilon.$
Taking the limit as $\epsilon\searrow 0$ yields the inequality
$\inf_{g(x)\in{C}}f(x)$$\leq$ $f(\overline{x})+c_{*}\,\mathrm{dist}\big{(}g(\overline{x}),{C}\big{)}.$
Since $\overline{x}$ was arbitrary in $\mathbb{R}^{n},$ it follows that
$\inf_{g(x)\in{C}}f(x)$$\leq$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c_{*}\,\mathrm{dist}\big{(}g(x),{C}\big{)}\big{\}}.$
Since the opposite inequality holds triviality, the desired equality follows. ∎
To study the exactness of penalty functions, we can imbed the problem (P0) into a parametrized family of mathematical programs. For each $u\in\mathbb{R}^{m}$ consider the perturbed problem
minimize $\textrm{minimize }f(x)\quad\textrm{ subject to }g(x)\in C+u.$
Accordingly, we define the optimal value function $V\colon\mathbb{R}^{m}\to\mathbb{R}\cup\{\pm\infty\}$ by
$V(u)$ ::=:= $\begin{cases}\inf_{g(x)\in C+u}f(x)&\textrm{ if }\{x\in\mathbb{R}^{n}\mid g(x)\in C+u\}\neq\emptyset,\\ \infty&\textrm{ otherwise.}\end{cases}$
Corollary 3.1.
If there exists a constant $c_{*}>0$ such that
$\inf_{g(x)\in{C}}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c_{*}\,\mathrm{dist}\big{(}g(x),{C}\big{)}\big{\}},$
then
$\liminf_{u\to 0}\frac{V(u)-V(0)}{\|u\|}$ $>$ -$\infty$.- ∞ .
The converse holds true if $f$ is bounded from below on $\mathbb{R}^{n}.$
Proof.
Assume that
$\inf_{g(x)\in{C}}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c_{*}\,\mathrm{dist}\big{(}g(x),{C}\big{)}\big{\}}.$
Take any $u\in\mathbb{R}^{m}.$ By Theorems 3.1 and 3.2, if there exists a point $\overline{x}\in\mathbb{R}^{n}$ with $g(\overline{x})\in{C}+u,$ then
$\inf_{g(x)\in{C}}f(x)$$\leq$ $f(\overline{x})+c_{*}\,\|u\|,$
which easily yields that
$V(0)$$\leq$ $V(u)+c_{*}\,\|u\|.$
This inequality is still valid when there is no point $\overline{x}$ with $g(\overline{x})\in{C}+u.$ Since $u\in\mathbb{R}^{m}$ was arbitrary, we get the desired conclusion
$\liminf_{u\to 0}\frac{V(u)-V(0)}{\|u\|}$$\geq$ $-c_{*}\ >\ -\infty.$
Now assume that $f$ is bounded from below on $\mathbb{R}^{n}$ and the converse is not true. By Theorem 3.2, there exists a sequence $(x_{k},u_{k})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ with $g(x_{k})\in C+u_{k}$ such that
$\inf_{g(x)\in{C}}f(x)$ $>$ $f(x_{k})+k\,\|u_{k}\|\ \geq\ \inf_{x\in\mathbb{R}^{n}}f(x)+k\,\|u_{k}\|.$
Certainly $u_{k}\neq 0$ and $u_{k}\to 0$ as $k\to\infty.$ Moreover, we have
00 0 $>$ $\big{[}f(x_{k})-\inf_{g(x)\in{C}}f(x)\big{]}+k\,\|u_{k}\|\ \geq\ V(u_{k})-V(0)+k\,\|u_{k}\|.$
Hence
00 0 $>$ $\frac{V(u_{k})-V(0)}{\|u_{k}\|}+k,$
which is impossible for $k$ sufficiently large. ∎
The converse of the corollary above is not true if $f$ is not bounded from below on $\mathbb{R}^{n};$ see Example 5.2(ii).
Remark 3.2.
The results presented in this section can be applied to an abstract constraint of the form $x\in X$ with $X$ being a subset of $\mathbb{R}^{n}$ by substituting $f$ by $f+\delta_{X}(\cdot).$
4. Exact penalty functions: the locally Lipschitz case
This section is inspired by the work of Zaslavski [37, 39]. Indeed, consider the problem (P) with the objective function $f$ being locally Lipschitz continuous and the constraint set $S$ being given by
$S$ ::=:= $\{x\in\mathbb{R}^{n}\ |\ g(x)\in C\},$
where $g\colon\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a locally Lipschitz mapping and $C$ is a nonempty closed subset of $\mathbb{R}^{m}.$ Assume that the optimal value $f_{*}:=\inf_{g(x)\in C}f(x)$ is finite. Under regularity conditions, local and at infinity, we will give a necessary and sufficient condition for the following to hold
$\inf_{g(x)\in C}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c\,\mathrm{dist}(g(x),C)\big{\}}$
for all $c$ sufficiently large.
Recall that for each vector $w\in\mathbb{R}^{m},$ the function $\langle w,g\rangle\colon\mathbb{R}^{n}\to\mathbb{R}$ is defined by
$\langle w,g\rangle(x)$ ::=:= $\langle w,g(x)\rangle\quad\textrm{ for }\quad x\in\mathbb{R}^{n}.$
Let us start with the following.
Definition 4.1.
We say that the (generalized) Mangasarian–Fromovitz constraint qualification (MFCQ) holds at $x\in S$ if there is no unit vector $w\in N_{C}(g(x))$ such that $0\in\partial\langle w,g\rangle(x).$
Remark 4.1.
When $C=\{0\}\times(\mathbb{R}_{-})^{m-p}\subset\mathbb{R}^{m}$ and $g$ is of class $C^{1},$ the above definition reduces to the traditional one.
Let $K_{\infty}(f,S)$ be the set of all $t\in\mathbb{R}$ for which there exists a sequence $x_{k}\in\mathbb{R}^{n}\setminus S$ satisfying the following conditions:
$\lim_{k\to\infty}\|x_{k}\|=\infty,\quad\lim_{k\to\infty}\|x_{k}\|\nu(x_{k})=0,\quad\lim_{k\to\infty}f(x_{k})=t,\quad\textrm{ and }\quad\lim_{k\to\infty}\mathrm{dist}(g(x_{k}),C)=0.$
Here and in the following we put
$\nu(x)$ ::=:= $\inf\|\lambda u+(1-\lambda)v\|\quad\textrm{ for }\quad x\in\mathbb{R}^{n}$
with the infimum being taken over all real numbers $\lambda\in(0,1)$ and vectors $u\in\partial f(x),v\in\partial\langle w,g\rangle(x)$ and $w\in N_{C}(g(x))$ with $\|w\|=1.$
Remark 4.2.
(i) The definition of $K_{\infty}(f,S)$ is strongly related to the weak Palais–Smale condition (see, for example, [24, 32]).
(ii) When $C=\{0\}\subset\mathbb{R}^{m}$ and $f$ and $g$ are of class $C^{1},$ we have that
$\nu(x)$ $=$ $\inf\big{\{}\|\lambda\nabla f(x)+(1-\lambda)\nabla\langle w,g\rangle(x)\|\mid\lambda\in(0,1),w\in\mathbb{R}^{m},\|w\|=1\big{\}}.$
(iii) Observe that if for some $c_{*}>0$ we have
$-\infty\ <\ \inf_{g(x)\in C}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c_{*}\,\mathrm{dist}(g(x),C)\big{\}},$
then the function
$\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{\infty\},\quad x\mapsto f(x)+c_{*}\,\mathrm{dist}(g(x),C),$
must be bounded from below. Conversely, we have the following result, which improves [8, Theorem 4.2] and [37, Theorems 1.5 and 3.1].
Theorem 4.1.
Let the following two conditions hold
  1. (C1)
    the (MFCQ) is satisfied at every global minimum of $f$ on $S;$
  2. (C2)
    the inclusion $K_{\infty}(f,S)\subset(f_{*},\infty)$ is valid.
Then the following properties are equivalent:
  1. (i)
    for all $c$ sufficiently large, one has
    $\inf_{g(x)\in C}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c\,\mathrm{dist}(g(x),C)\big{\}};$
  2. (ii)
    there exists a constant $c_{*}>0$ such that
    $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c_{*}\mathrm{dist}(g(x),C)\big{\}}$ $>$ -$\infty$.- ∞ .
When these equivalent properties hold, one has for any $c>c_{*}$ that
$\mathrm{argmin}_{g(x)\in C}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big{\{}f(x)+c\,\mathrm{dist}(g(x),C)\big{\}};$
moreover, if $\overline{x}$ is a global minimizer for the problem (P0), then there exists a vector $w\in N_{C}(g(\overline{x}))\cap\mathbb{B}$ satisfying the following optimality condition
00 0 $\in$ $\partial f(\overline{x})+c\,\partial\langle w,g\rangle(\overline{x}).$
To prove the theorem, we need the following simple fact.
Lemma 4.1.
Let $x_{k},v_{k}\in\mathbb{R}^{n}$ and $w_{k}\in\mathbb{R}^{m}$ be sequences converging to $\overline{x},\overline{v}$ and $\overline{w},$ respectively, such that $v_{k}\in\partial\langle w_{k},g\rangle(x_{k})$ for all $k.$ Then $\overline{v}\in\partial\langle\overline{w},g\rangle(\overline{x}).$
Proof.
Since $v_{k}\in\partial\langle w_{k},g\rangle(x_{k}),$ it follows from Lemma 2.8 that $v_{k}\in D^{*}g(x_{k})(w_{k}),$ which by definition is equivalent to $(v_{k},-w_{k})\in N_{\mathrm{gph}g}(x_{k},g(x_{k})).$ Letting $k\to\infty$ and applying Lemma 2.1, we get $(\overline{v},-\overline{w})\in N_{\mathrm{gph}g}(\overline{x},g(\overline{x})),$ and so $\overline{v}\in D^{*}g(\overline{x})(\overline{w}).$ By Lemma 2.8 again, $\overline{v}\in\partial\langle\overline{w},g\rangle(\overline{x}),$ as required. ∎
Proof of Theorem 4.1.
We first prove the equivalence between (i) and (ii). To see this, define the function $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ by
$\psi(x)$ ::=:= $\mathrm{dist}(g(x),C)\quad\textrm{ for }\quad x\in\mathbb{R}^{n},$
which is locally Lipschitz continuous and residual for $S.$ By Theorem 3.1, it suffices to show the equivalence of (ii) with the following property:
  1. ($\star$)
    there exists a constant $c_{*}>0$ such that
    $c_{*}\psi(x)$$\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad x\in\mathbb{R}^{n}.$
Indeed, the ($\star$) $\Rightarrow$ (ii) part is trivial. For the (ii) $\Rightarrow$ ($\star$) part we likewise proceed by contradiction. Let $c_{*}>0$ be such that
$\inf_{x\in\mathbb{R}^{n}}\left\{f(x)+c_{*}\psi(x)\right\}$ $>$ -$\infty$,- ∞ ,
and suppose that for each $k>c_{*},$ there exists $x_{k}\in\mathbb{R}^{n}$ such that
$k\psi(x_{k})$ $<$ $[f_{*}-f(x_{k})]_{+}.$
Since the function $\psi$ is nonnegative, $f_{*}-f(x_{k})>0.$ Moreover, if $\psi(x_{k})=0$ then $x_{k}\in S$ and so $f_{*}\leq f(x_{k}),$ a contradiction. Therefore
00 0 $<$ $k\psi(x_{k})\ <\ f_{*}-f(x_{k}).$ (2)
Define the function $F_{k}\colon\mathbb{R}^{n}\to\mathbb{R},x\mapsto F_{k}(x),$ by
$F_{k}(x)$ ::=:= $f(x)+k\psi(x),$
which is locally Lipschitz continuous and satisfies
$-\infty\ <\ \mathfrak{m}:=\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c_{*}\psi(x)\big{\}}$$\leq$ $\inf_{x\in\mathbb{R}^{n}}F_{k}(x)\ <\ f_{*},$
where the last inequality follows from (2).
Put
$\varepsilon_{k}:=f_{*}-\inf_{x\in\mathbb{R}^{n}}F_{k}(x)\quad\textrm{ and }\quad\lambda_{k}:=\tfrac{\|x_{k}\|+1}{2}.$
Then $0<\varepsilon_{k}<f_{*}-\mathfrak{m}$ and we deduce from (2) that
$F_{k}(x_{k})$ $<$ $f_{*}\ =\ \inf_{x\in\mathbb{R}^{n}}F_{k}(x)+\varepsilon_{k}.$
By the Ekeland variational principle (see [14]), there exists $x_{k}^{\prime}\in\mathbb{R}^{n}$ having the following properties:
  1. (a)
    $F_{k}(x_{k}^{\prime})\leq F_{k}(x_{k})<f_{*}=\inf_{x\in\mathbb{R}^{n}}F_{k}(x)+\varepsilon_{k},$
  2. (b)
    $\|x_{k}^{\prime}-x_{k}\|\leq\lambda_{k},$ and
  3. (c)
    $F_{k}(x_{k}^{\prime})\leq F_{k}(x)+\frac{\varepsilon_{k}}{\lambda_{k}}\|x-x_{k}^{\prime}\|$ for all $x\in\mathbb{R}^{n}$.
Note that $\psi(x_{k}^{\prime})>0$ because otherwise we would have $x_{k}^{\prime}\in S,$ so that
$F_{k}(x_{k}^{\prime})$ $=$ $f(x_{k}^{\prime})+k\psi(x_{k}^{\prime})\ =\ f(x_{k}^{\prime})\ \geq\ f_{*}$
in contradiction to the property (a). Moreover, we have
$\mathfrak{m}$$\leq$ $f(x_{k}^{\prime})+{c_{*}}\psi(x_{k}^{\prime})\ \leq\ f(x_{k}^{\prime})+k\psi(x_{k}^{\prime})\ <\ f_{*}.$ (3)
It follows that
$\mathfrak{m}+(k-c_{*})\psi(x_{k}^{\prime})$$\leq$ $f_{*}.$
Consequently, $\psi(x_{k}^{\prime})$ tends to 00 as $k\to\infty.$ This when combined with (3) implies that the sequence $f(x_{k}^{\prime})$ is bounded and has all its cluster points in the interval $[\mathfrak{m},f_{*}].$
From the property (b), we easily deduce that
$\frac{\|x_{k}\|-1}{2}$$\leq$ $\|x_{k}^{\prime}\|\ \leq\ \frac{3\|x_{k}\|+1}{2}.$ (4)
In particular, the sequence $x_{k}$ is bounded if and only if the sequence $x_{k}^{\prime}$ is bounded.
On the other hand, by the property (c), $x_{k}^{\prime}$ is a minimizer of the locally Lipschitz function
$\mathbb{R}^{n}\to\mathbb{R},\quad x\mapsto F_{k}(x)+\frac{\varepsilon_{k}}{\lambda_{k}}\|x-x_{k}^{\prime}\|.$
Lemma 2.5 (Fermat rule) when combined with Lemma 2.4 (sum rule) yields
00 0 $\in$ $\partial\left[F_{k}+\frac{\varepsilon_{k}}{\lambda_{k}}\|\cdot-x_{k}^{\prime}\|\right](x_{k}^{\prime})$
$\subset$ $\partial f(x_{k}^{\prime})+k\partial\psi(x_{k}^{\prime})+\frac{\varepsilon_{k}}{\lambda_{k}}\mathbb{B}.$
Note that $g(x_{k}^{\prime})\not\in C$ because $\psi(x_{k}^{\prime})>0.$ Moreover, we know from Lemmas 2.2 and 2.7 that
$\partial\psi(x_{k}^{\prime})$$\subset$ $\bigcup_{w}\partial\langle w,g\rangle(x_{k}^{\prime}),$
where the union is taken on all vectors $w$ belonging to the set
$\frac{g(x_{k}^{\prime})-\Pi_{C}(g(x_{k}^{\prime}))}{\psi(x_{k}^{\prime})},$
which by definition is a subset of the limiting normal cone $N_{C}(g(x_{k}^{\prime})).$ Therefore, we can find vectors $u_{k}\in\partial f(x_{k}^{\prime})$ and $v_{k}\in\partial\langle w_{k},g\rangle(x_{k}^{\prime})$ for some $w_{k}\in N_{C}(g(x_{k}^{\prime}))$ with $\|w_{k}\|=1$ such that
00 0 $\in$ $u_{k}+kv_{k}+\frac{\varepsilon_{k}}{\lambda_{k}}\mathbb{B}.$ (5)
There are two cases to be considered.
Case 1: the sequence $\{x_{k}\}$ is bounded.
By (4), the sequence $\{x_{k}^{\prime}\}$ is bounded too. Passing to a subsequence if necessary, we can suppose that $x_{k}^{\prime}$ converges to some point $\overline{x}\in\mathbb{R}^{n}.$ Certainly, $\psi(\overline{x})=0$ (equivalently, $\overline{x}\in S$) and $f(\overline{x})=f_{*},$ which imply that $\overline{x}$ is a global minimizer of $f$ on $S.$
On the other hand, since the mappings $f$ and $g$ are Lipschitz continuous around $\overline{x}$ and $\|w_{k}\|=1,$ it follows from Lemma 2.6 that the sequences $u_{k}$ and $v_{k}$ must be bounded. So we may assume that the sequences $u_{k},v_{k}$ and $w_{k}$ converge to vectors $\overline{u},\overline{v}$ and $\overline{w},$ respectively. Certainly, $\overline{u}\in\partial f(\overline{x}),$ $\overline{v}\in\partial\langle\overline{w},g\rangle(\overline{x})$ and $\overline{w}\in N_{C}(g(\overline{x}))$ with $\|\overline{w}\|=1$ (due to Lemmas 2.1, 2.3 and 4.1).
Observe that (5) can be rewritten as
00 0 $\in$ $\frac{1}{k}u_{k}+v_{k}+\frac{1}{k}\frac{\varepsilon_{k}}{\lambda_{k}}\mathbb{B}.$
Equivalently,
$\left\|\frac{1}{k}u_{k}+v_{k}\right\|$$\leq$ $\frac{1}{k}\frac{\varepsilon_{k}}{\lambda_{k}}\ <\ \frac{2(f_{*}-\mathfrak{m})}{k(\|x_{k}\|+1)}.$
Letting $k\to\infty,$ one gets $\overline{v}=0,$ in contradiction to (C1).
Case 2: the sequence $\{x_{k}\}$ is unbounded.
By (4), the sequence $\{x_{k}^{\prime}\}$ is also unbounded. Passing to a subsequence if necessary, we can suppose that $\|x_{k}^{\prime}\|\to\infty$ as $k\to\infty.$ Note that (5) can be rewritten as
00 0 $\in$ $\frac{1}{{1+k}}u_{k}+\frac{k}{{1+k}}v_{k}+\frac{1}{{1+k}}\cdot\frac{\varepsilon_{k}}{\lambda_{k}}\mathbb{B}.$
By definition, then
$\nu(x_{k}^{\prime})$$\leq$ $\left\|\frac{1}{{1+k}}u_{k}+\frac{k}{{1+k}}v_{k}\right\|\ \leq\ \frac{\varepsilon_{k}}{(1+k)\lambda_{k}}\ <\ \frac{2(f_{*}-\mathfrak{m})}{{(1+k)}(\|x_{k}\|+1)}.$
In combining this with (4), we obtain
$(\|x_{k}^{\prime}\|+1)\nu(x_{k}^{\prime})$ $<$ $\frac{(f_{*}-\mathfrak{m})}{3({1+k})}.$
Thus, $\|x_{k}^{\prime}\|\nu(x_{k}^{\prime})\to 0$ as $k\to\infty.$ On the other hand, we have shown that $x_{k}^{\prime}\not\in S,$ $\psi(x_{k}^{\prime})\to 0$ as $k\to\infty$ and the sequence $f(x_{k}^{\prime})$ has a cluster point, say, $t,$ which belongs to the interval $[\mathfrak{m},f_{*}].$ Certainly $t\in K_{\infty}(f,S),$ in contradiction to (C2). The equivalence of (ii) with ($\star$), and hence with (i), is proved.
Finally, assume that the property (i) holds. By Theorem 3.1, one has for any $c>c_{*}$ that
$\mathrm{argmin}_{g(x)\in C}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big{\{}f(x)+c\,\mathrm{dist}(g(x),C)\big{\}}.$
Let $\overline{x}$ be a global minimum for the problem (P0). Then $\overline{x}$ is also a global minimum for the unconstrained optimization problem
$\textrm{minimize}_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c\,\mathrm{dist}(g(x),C)\big{\}},$
and so $0\in\partial\big{\{}f(\cdot)+c\,\mathrm{dist}(g(\cdot),C)\big{\}}(\overline{x})$ due to Lemma 2.5. This, together with Lemmas 2.2, 2.4 and 2.7, yields
00 0 $\in$ $\partial f(\overline{x})+c\,\partial\langle w,g\rangle(\overline{x})$
for some $w\in N_{C}(g(\overline{x}))\cap\mathbb{B},$ as required. ∎
The conditions (C1) and (C2) cannot be omitted as shown below.
Example 4.1.
(i) Let $n:=1,S:=\{x\in\mathbb{R}\mid x^{2}=0\}$ and $f(x):=x^{3}.$ Clearly, for any $c>0$ we have
$\inf_{x\in S}f(x)=0>-\infty=\inf_{x\in\mathbb{R}}\big{\{}f(x)+c\,\psi(x)\big{\}},$
where $\psi(x):=x^{2}.$ Observe that the condition (C2) holds while (C1) does not.
(ii) Let $n:=2,S:=\{x:=(x_{1},x_{2})\in\mathbb{R}^{2}\ |\ x_{1}=0\}$ and $f(x):=e^{x_{1}x_{2}}.$ Clearly, for any $c>0$ we have
$\inf_{x\in S}f(x)=1>0=\inf_{x\in\mathbb{R}^{2}}\big{\{}f(x)+c\,\psi(x)\big{\}},$
where $\psi(x):=|x_{1}|.$ Observe that the condition (C1) is satisfied, but the condition (C2) is not.
(iii) Let $n:=2,S:=\{x:=(x_{1},x_{2})\in\mathbb{R}^{2}\mid x_{1}-x_{2}=0\}$ and $f(x):=x_{1}-x_{2}.$ Observe that the assumption (A1) in [8], the assumption (1.1) in [37] and the assumption (1.6) in [39] do not satisfy, and so we can not apply the results in these papers. On the other hand, it is easy to see that the conditions (C1) and (C2) hold, and indeed, for any $c\geq 1$ we have
$\inf_{x\in S}f(x)=0=\inf_{x\in\mathbb{R}^{2}}\big{\{}f(x)+c\,\psi(x)\big{\}},$
where $\psi(x):=|x_{1}-x_{2}|.$
(iv) Let $n:=2,S:=\{x:=(x_{1},x_{2})\in\mathbb{R}^{2}\mid x_{1}=0\}$ and $f(x):=x_{1}^{3}+x_{2}^{2}.$ For any $c>0$ we have
$\inf_{x\in S}f(x)=0>-\infty=\inf_{x\in\mathbb{R}^{2}}\big{\{}f(x)+c\,\psi(x)\big{\}},$
where $\psi(x):=|x_{1}|.$ Observe that the conditions (C1) and (C2) are satisfied.
Remark 4.3.
(i) Theorem 4.1 can be strengthened by replacing $f$ by $f_{0}+\delta_{X}$ with $f_{0}\colon\mathbb{R}^{n}\rightarrow\mathbb{R}$ being a locally Lipschitz function and $X$ being a closed subset of $\mathbb{R}^{n}.$ Since we do not use this fact, we leave the details for the interested reader.
(ii) The condition (C2) holds trivially when $K_{\infty}(f,S)$ is an empty set. Consequently, Theorem 4.1 improves [8, Theorem 4.2] where it is assumed that $C=\mathbb{R}_{-}^{m},$ $f$ and $g$ are twice continuously differentiable functions, $f$ is radially unbounded on a perturbation of the constraint set $S$ and the linear independence constraint qualification holds at every point in $S.$ Theorem 4.1 also improves [37, Theorems 1.5 and 3.1] where it is assumed that $C=\{0\}\times\mathbb{R}_{-}^{m-p},$ $f$ and $g$ are locally Lipschitz, $f$ is coercive and the linear independence constraint qualification holds at every global minimum of $f$ on $S.$
5. Exact penalty functions: the semi-algebraic case
In this section we investigate exact penalty functions for constrained optimization problems with semi-algebraic data. The main tool for this study will be Łojasiewicz inequalities on unbounded sets.
5.1. Łojasiewicz inequalities on unbounded sets
In this subsection, we present some Łojasiewicz inequalities on unbounded sets for (not necessarily continuous) semi-algebraic functions, and for the convenience of the reader, we provide detailed proofs to make the paper self-contained. The results presented here, which are inspired by the work of Dinh, Hà and Thao [12] (see also [11, 23, 25, 26]), will play a crucial role in establishing exact penalty properties in constrained optimization with semi-algebraic data.
Let us start with the following definition.
Definition 5.1.
Let $X$ be a subset of $\mathbb{R}^{n}$ and $\phi,\psi\colon X\rightarrow\mathbb{R}$ be functions. A sequence $\{x_{k}\}$ in $X$ with $\|x_{k}\|\to\infty$ is said to be
  1. (i)
    a sequence of the first type for the pair $(\phi,\psi)$ if $\psi(x_{k})\to 0$ and $|\phi(x_{k})|\geq\epsilon$ for some $\epsilon>0;$
  2. (ii)
    a sequence of the second type for the pair $(\phi,\psi)$ if the sequence $\psi(x_{k})$ is bounded and $|\phi(x_{k})|\to\infty.$
The following result improves [12, Proposition 3.7] in which $X=\mathbb{R}^{n},$ $\psi$ is the absolute value function of a polynomial and $\phi$ is the distance function $x\mapsto\mathrm{dist}(x,\psi^{-1}(0)).$
Lemma 5.1.
Let $X$ be a semi-algebraic subset of $\mathbb{R}^{n}$ and $\phi,\psi\colon X\rightarrow\mathbb{R}$ be nonnegative and semi-algebraic functions. The following two conditions are equivalent:
  1. (i)
    there exist some constants $c>0,\epsilon>0$ and $\alpha\in(0,1]$ such that
    $c\,[\psi(x)]^{\alpha}$$\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in\psi^{-1}([0,\epsilon));$
  2. (ii)
    there is no sequence of the first type for the pair $(\phi,\psi).$
Proof.
(i) $\Rightarrow$ (ii) The implication is straightforward.
(ii) $\Rightarrow$ (i) Since there is no sequence of the first type for the pair $(\phi,\psi),$ we have
  1. Fact 1.
    $\phi(x)\to 0$ as $\psi(x)\to 0;$ in particular, $\phi(x)=0$ provided that $\psi(x)=0.$
On the other hand, in light of Theorem 2.1, $\psi(X)$ is a semi-algebraic set in $\mathbb{R},$ so it is a finite union of points and intervals (see Lemma 2.9). Consequently, there exists a constant $\epsilon>0$ such that either $(0,\epsilon)\cap\psi(X)=\emptyset$ or $(0,\epsilon)\subset\psi(X).$ In the first case, for all $x\in\psi^{-1}[0,\epsilon)$ we have $\psi(x)=0,$ and by Fact 1, $\phi(x)=0;$ consequently, (ii) holds trivially for any choice of $c>0$ and $\alpha>0.$ So, assume that the second case holds, i.e., $(0,\epsilon)\subset\psi(X).$ Then for all $t\in(0,\epsilon),$ the set $\psi^{-1}(t)$ is nonempty and so the function
$\mu\colon(0,\epsilon)\to\mathbb{R}\cup\{\infty\},\quad t\mapsto\sup_{x\in\psi^{-1}(t)}\phi(x),$
is well defined. Observe that, by Theorem 2.1, the function $\mu$ is semi-algebraic. Moreover, it follows from Fact 1 that $\mu(t)\to 0$ as $t\to 0^{+}.$ By Lemma 2.11 and decreasing $\epsilon$ if necessary, we can see that the function $\mu$ is finite and either constant or strictly monotone. There are two cases to consider.
Case 1.1:
$\mu$ is constant.
Then $\mu(t)=0$ for all $t\in(0,\epsilon).$ It follows that for all $x\in\psi^{-1}(0,\epsilon),$
$\psi(x)\geq 0=\phi(x).$
Then (ii) holds trivially for any choice of $c>0$ and $\alpha>0.$
Case 1.2:
$\mu$ is not constant.
Thanks to Lemma 2.10(i), we can write
$\mu(t)=a\,t^{\alpha}+o(t^{\alpha})\quad\textrm{ as }\quad t\to 0^{+}$
for some constants $a>0$ and $\alpha\in\mathbb{Q}.$ Observe that $\alpha>0$ because $\mu(t)\to 0$ as $t\to 0^{+}.$ Reducing $\epsilon$ and replacing $\alpha$ by $\min\{\alpha,1\}$ (if necessary), we can see that
$\mu(t)\leq 2a\,t^{\alpha}\quad\textrm{ for }\quad t\in(0,\epsilon).$
Hence (ii) holds when $c:=2a>0.$
The following corollary improves [25, Lemma 3.2] in which $X$ and $\phi$ are assumed to be compact and continuous, respectively; see also [23, Theorem 2.10].
Corollary 5.1.
Let $X$ be a semi-algebraic subset of $\mathbb{R}^{n}$ and $\phi,\psi\colon X\rightarrow\mathbb{R}$ be nonnegative and semi-algebraic functions. If in addition $\phi$ is bounded from above, then the following two conditions are equivalent:
  1. (i)
    there exist some constants $c>0$ and $\alpha\in(0,1]$ such that
    $c\,[\psi(x)]^{\alpha}$$\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in X;$
  2. (ii)
    there is no sequence of the first type for the pair $(\phi,\psi).$
Proof.
(i) $\Rightarrow$ (ii) The implication is straightforward.
(ii) $\Rightarrow$ (i) By Lemma 5.1, we can find constants $c_{1}>0,\epsilon>0$ and $\alpha\in(0,1]$ such that
$c_{1}[\psi(x)]^{\alpha}$$\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in\psi^{-1}([0,\epsilon)).$
On the other hand, since $\phi$ is bounded from above, there exists a constant $M>0$ such that $\phi(x)\leq M$ for all $x\in X.$ Hence for all $x\in\psi^{-1}([\epsilon,\infty))$ we have
$[\psi(x)]^{\alpha}$$\geq$ $\epsilon^{\alpha}\ =\ \frac{\epsilon^{\alpha}}{M}M\ \geq\ \frac{\epsilon^{\alpha}}{M}\,\phi(x).$
Letting $c:=\max\{c_{1},\frac{M}{\epsilon^{\alpha}}\},$ we get the desired conclusion. ∎
Note here that the assumption of boundedness of the function $\phi$ cannot be omitted; see Example 5.1.
Corollary 5.2.
Let $A,B$ be semi-algebraic subsets of $\mathbb{R}^{n}$ such that $A\cap B\neq\emptyset.$ Then there exist some constants $c>0$ and $\alpha\in(0,1]$ such that
$c\,\big{(}\mathrm{dist}(x,A)+\mathrm{dist}(x,B)\big{)}^{\alpha}$$\geq$ $\frac{\mathrm{dist}(x,A\cap B)}{1+\|x\|^{2}}\quad\textrm{ for all}\quad x\in\mathbb{R}^{n}.$
Proof.
Define the functions $\phi,\psi\colon\mathbb{R}^{n}\to\mathbb{R}$ by
$\phi(x):=\frac{\mathrm{dist}(x,A\cap B)}{1+\|x\|^{2}}\quad$ and $\quad\psi(x):=\mathrm{dist}(x,A)+\mathrm{dist}(x,B),$
which are nonnegative, continuous and semi-algebraic (by Theorem 2.1). Note that $\phi(x)\to 0$ as $\|x\|\to\infty.$ In particular, the function $\phi$ is bounded from above and there is no sequence of the first type for the pair $(\phi,\psi).$ The desired conclusion follows directly from Corollary 5.1. ∎
Lemma 5.2.
Let $X$ be a semi-algebraic subset of $\mathbb{R}^{n}$ and $\phi,\psi\colon X\rightarrow\mathbb{R}$ be nonnegative and semi-algebraic functions. The following two conditions are equivalent:
  1. (i)
    there exist some constants $c>0,R>0$ and $\beta\geq 1$ such that
    $c\,[\psi(x)]^{\beta}$$\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in\psi^{-1}((R,\infty));$
  2. (ii)
    there is no sequence of the second type for the pair $(\phi,\psi).$
Proof.
(i) $\Rightarrow$ (ii) The implication is straightforward.
(ii) $\Rightarrow$ (i) Since there is no sequence of the second type, we have
  1. Fact 2.
    For any $a,b\in\mathbb{R}$ with $0\leq a\leq b<\infty,$ there exists a constant $M>0$ such that $\phi(x)\leq M$ for all $x\in\psi^{-1}([a,b]).$
On the other hand, in light of Theorem 2.1, $\psi(X)$ is a semi-algebraic set in $\mathbb{R},$ so it is a finite union of points and intervals (see Lemma 2.9). Consequently, there exists a constant $R>0$ such that either $(R,\infty)\cap\psi(X)=\emptyset$ or $(R,\infty)\subset\psi(X).$ If the first case happens, then (ii) holds trivially for any choice of $c>0$ and $\beta>0$ and there is nothing to prove. Hence, assume that $(R,\infty)\subset\psi(X).$ Then for all $t\in(R,\infty),$ the set $\psi^{-1}(t)$ is nonempty. It follows from Fact 2 that the function
$\mu\colon(R,\infty)\to\mathbb{R},\quad t\mapsto\sup_{x\in\psi^{-1}(t)}\phi(x),$
is well defined. Observe that, by Theorem 2.1, the function $\mu$ is semi-algebraic. By Lemma 2.11 and increasing $R$ if necessary, we can see that the function $\mu$ is either constant or strictly monotone. There are two cases to consider.
Case 2.1:
$\mu$ is constant, say, $c^{\prime}.$
Then $c^{\prime}\geq 0$ and for all $x\in\psi^{-1}(R,\infty)$ we have
$\psi(x)\geq R=\frac{R}{c^{\prime}+1}(c^{\prime}+1)\geq\frac{R}{c^{\prime}+1}\phi(x).$
Then (ii) holds trivially when $c:=\frac{c^{\prime}+1}{R}$ and $\beta:=1.$
Case 2.2:
$\mu$ is not constant.
Thanks to Lemma 2.10(ii), we can write
$\mu(t)=b\,t^{\beta}+o(t^{\beta})\quad\textrm{ as }\quad t\to\infty$
for some constants $b>0$ and $\beta\in\mathbb{Q}.$ Increasing $R$ and replacing $\beta$ by $\max\{\beta,1\}$ (if necessary), we can see that
$\mu(t)\leq 2b\,t^{\beta}\quad\textrm{ for }\quad t\in(R,\infty).$
Hence (ii) holds when $c:=2b>0.$
The following result improves [12, Proposition 3.10] in which $X=\mathbb{R}^{n},$ $\psi$ is the absolute value function of a polynomial and $\phi$ is the distance function $x\mapsto\mathrm{dist}(x,\psi^{-1}(0)).$
Proposition 5.1.
Let $X$ be a semi-algebraic subset of $\mathbb{R}^{n}$ and $\phi,\psi\colon X\rightarrow\mathbb{R}$ be nonnegative and semi-algebraic functions. The following two conditions are equivalent:
  1. (i)
    there exist some constants $c>0,\alpha>0,$ and $\beta>0$ with $\alpha\leq 1\leq\beta$ such that
    $c\left([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\right)$$\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in X;$
  2. (ii)
    there are no sequences of the first and second types for the pair $(\phi,\psi).$
Proof.
(i) $\Rightarrow$ (ii) The implication is straightforward.
(ii) $\Rightarrow$ (i) By Lemma 5.1, there exist some constants $c_{1}>0,\epsilon>0$ and $\alpha\in(0,1]$ such that
$c_{1}\,[\psi(x)]^{\alpha}$$\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in\psi^{-1}([0,\epsilon)).$
In view of Lemma 5.2, there exist some constants $c_{2}>0,R>\epsilon$ and $\beta\geq 1$ such that
$c_{2}\,[\psi(x)]^{\beta}$$\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in\psi^{-1}((R,\infty)).$
On the other hand, it follows from Fact 2 in the proof of Lemma 5.2 that there exists a constant $M>0$ such that for all $x\in\psi^{-1}([\epsilon,R])$ we have $\phi(x)\leq M,$ and hence
$[\psi(x)]^{\alpha}+[\psi(x)]^{\beta}$$\geq$ $\epsilon^{\alpha}+\epsilon^{\beta}\ =\ \frac{\epsilon^{\alpha}+\epsilon^{\beta}}{M}M\ \geq\ \frac{\epsilon^{\alpha}+\epsilon^{\beta}}{M}\,\phi(x).$
Letting $c:=\max\{c_{1},c_{2},\frac{M}{\epsilon^{\alpha}+\epsilon^{\beta}}\},$ we get the desired conclusion. ∎
The following simple example shows that the exponents $\alpha$ and $\beta$ in Proposition 5.1 are different in general.
Example 5.1.
Consider the polynomial functions $\phi(x_{1},x_{2}):=x_{1}^{2}+x_{2}^{2}$ and $\psi(x_{1},x_{2}):=x_{1}^{2}+x_{2}^{4}$ on $\mathbb{R}^{2}.$ It is not hard to see that there are no constants $c>0$ and $\alpha>0$ such that
$c\,[\psi(x_{1},x_{2})]^{\alpha}$$\geq$ $\phi(x_{1},x_{2})\quad\textrm{ for all }\quad(x_{1},x_{2})\in\mathbb{R}^{2}.$
On the other hand, it holds that
$[\psi(x_{1},x_{2})]^{\frac{1}{2}}+\psi(x_{1},x_{2})$$\geq$ $\phi(x_{1},x_{2})\quad\textrm{ for all }\quad(x_{1},x_{2})\in\mathbb{R}^{2}.$
5.2. Necessary and sufficient conditions for exact penalization
The results presented in this subsection are inspired by the works of Warga [34, Theorem 1], Dedieu [10, Theorem 3.1] and Luo, Pang and Ralph [27, Chapter 2]). Indeed, consider the problem (P) with $S$ being a nonempty semi-algebraic subset of $\mathbb{R}^{n}$ and $f\colon\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ being a semi-algebraic function. Assume that the optimal value $f_{*}:=\inf_{S}f$ is finite and let $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ be a residual function for $S$ which is semi-algebraic.
Remark 5.1.
If for some $c_{*}>0$ we have
$\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c_{*}\,\psi(x)\big{\}},$
then by Theorem 3.1, for any sequence of points $x_{k}$ in $\mathbb{R}^{n}$ with $\psi(x_{k})\to 0$ the sequence $[f_{*}-f(x_{k})]_{+}$ must converge to $0$ The converse does not hold in general; see Example 5.2(ii) below. On the other hand, we have the following result.
Theorem 5.1.
Assume that for any sequence of points $x_{k}$ in $\mathbb{R}^{n}$ with $\psi(x_{k})\to 0$ one has $[f_{*}-f(x_{k})]_{+}\to 0.$ Then there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that for all $c>c_{*},$
$\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c\,\big{(}1+[f(x)]^{2}\big{)}[\psi(x)]^{\alpha}\big{\}},$
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big{\{}f(x)+c\,\big{(}1+[f(x)]^{2}\big{)}[\psi(x)]^{\alpha}\big{\}}.$
Proof.
Let $A:=(-\infty,f_{*}]\times\{0\}$ and $B:=\mathrm{cl}(f,\psi)(\mathrm{dom}f).$ Then $A$ and $B$ are closed semi-algebraic sets in $\mathbb{R}^{2}.$ Moreover, our assumption yields $A\cap B=\{(f_{*},0)\}.$ In view of Corollary 5.2, there are some constants $c_{*}>0$ and $\alpha\in(0,1]$ such that
$c_{*}\big{(}\mathrm{dist}(y,A)+\mathrm{dist}(y,B)\big{)}^{\alpha}$$\geq$ $\frac{\mathrm{dist}(y,A\cap B)}{1+\|y\|^{2}}\quad\textrm{ for all}\quad y\in\mathbb{R}^{2}.$
Take $x\in\mathbb{R}^{n}$ with $f(x)<f_{*}$ and let $y:=(f(x),0).$ We have
$\mathrm{dist}(y,A)$ $=$ $0,$
$\mathrm{dist}(y,B)$$\leq$ $\|(f(x),0)-(f(x),\psi(x))\|=\psi(x),$
$\mathrm{dist}(y,A\cap B)$ $=$ $\|(f(x),0)-(f_{*},0)\|=f_{*}-f(x).$
Hence
$c_{*}[\psi(x)]^{\alpha}$$\geq$ $\frac{f_{*}-f(x)}{1+[f(x)]^{2}},$
or equivalently,
$c_{*}{\big{(}1+[f(x)]^{2}\big{)}}[\psi(x)]^{\alpha}$$\geq$ ${f_{*}-f(x)}.$
Clearly, this inequality is true for any $x\in\mathbb{R}^{n}$ with $f(x)\geq f_{*}.$ By Theorem 3.1, we get the desired conclusion. ∎
The following example shows that the converse of Theorem 5.1 does not hold and the term $(1+[f(x)]^{2})$ cannot be omitted.
Example 5.2.
(i) Let $n:=1,S:=\{0\}$ and consider the semi-algebraic functions
$f(x)$ ::=:= $x\quad\textrm{ and }\quad\psi(x)\ :=\ \frac{|x|}{1+x^{2}}.$
Then $\psi$ is a residual function for $S$ and it holds that
$\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}}\big{\{}f(x)+(1+[f(x)]^{2})[\psi(x)]\big{\}},$
while
$\psi(x)\to 0\quad\textrm{ and }\quad[f_{*}-f(x)]_{+}\to\infty\quad\textrm{ as }\ x\to-\infty.$
(ii) Let $n:=1,S:=\{0\}$ and $f(x):=x.$ Consider the semi-algebraic function $\psi\colon\mathbb{R}\to\mathbb{R}_{+}$ defined by
$\psi(x)$ ::=:= $\begin{cases}|x|&\textrm{ if }|x|\leq 1,\\ 1&\textrm{ otherwise,}\end{cases}$
which is a residual function for $S.$ A direct calculation shows that
$\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}}\big{\{}f(x)+(1+[f(x)]^{2})[\psi(x)]\big{\}},$
while for any $c>0$ and any $\alpha>0,$
$\inf_{x\in S}f(x)$ $=$ $0\ >\ -\infty\ =\ \inf_{x\in\mathbb{R}}\big{\{}f(x)+c\,[\psi(x)]^{\alpha}\big{\}}.$
The following result shows that the term $(1+[f(x)]^{2})$ can be removed provided that the function $f$ is bounded from below on $\mathbb{R}^{n}.$
Theorem 5.2.
If $f$ is bounded from below on $\mathbb{R}^{n},$ then the following properties are equivalent:
  1. (i)
    there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that for all $c>c_{*},$
    $\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c\,[\psi(x)]^{\alpha}\big{\}};$
  2. (ii)
    there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that
    $c_{*}[\psi(x)]^{\alpha}$$\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad x\in\mathbb{R}^{n};$
  3. (iii)
    for any sequence of points $x_{k}$ in $\mathbb{R}^{n}$ with $\psi(x_{k})\to 0,$ one has $[f_{*}-f(x_{k})]_{+}\to 0.$
When these equivalent properties hold, one has moreover for all $c>c_{*}$ that
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big{\{}f(x)+c\,[\psi(x)]^{\alpha}\big{\}}.$
Proof.
The equivalence between (i) and (ii) as well as the last conclusion follow from Theorem 3.1. To get the equivalence of (ii) and (iii), we can apply Corollary 5.1 to the functions $\phi$ and $\psi$ with $\phi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ being defined by
$\phi(x)$ ::=:= $[f_{*}-f(x)]_{+}\quad\textrm{ for }\quad x\in\mathbb{R}^{n},$
which is bounded from above on $\mathbb{R}^{n}.$
Remark 5.2.
(i) The assumption in Theorem 5.2 that $f$ is bounded from below on $\mathbb{R}^{n}$ cannot be dropped; see Example 5.2(ii). On the other hand, we have Theorem 5.3 below.
(ii) A semi-algebraic version of [34, Theorem 1] is followed directly from Theorem 5.2. To see this, let $f:=f_{0}+\delta_{X},$ $f_{0}\colon X\to\mathbb{R}$ be continuous and semi-algebraic, $X\subset\mathbb{R}^{n}$ and $S\subset X$ nonempty, compact and semi-algebraic. Also, let $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ be a residual function for $S$ which is continuous and semi-algebraic. Then it is not hard to see that $f$ is bounded from below on $\mathbb{R}^{n}$ and that the condition (iii) of Theorem 5.2 is satisfied, and so there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that for all $c>c_{*},$
$\inf_{x\in S}f_{0}(x)$ $=$ $\inf_{x\in X}\big{\{}f_{0}(x)+c\,[\psi(x)]^{\alpha}\big{\}},$
$\mathrm{argmin}_{S}f_{0}(x)$ $=$ $\mathrm{argmin}_{X}\big{\{}f_{0}(x)+c\,[\psi(x)]^{\alpha}\big{\}}.$
Theorem 5.3.
The following properties are equivalent:
  1. (i)
    there exist constants $c_{*}>0,\alpha>0,$ and $\beta>0$ with $\alpha\leq 1\leq\beta$ such that for all $c>c_{*},$
    $\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big{\{}f(x)+c\left([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\right)\big{\}};$
  2. (ii)
    there exist constants $c_{*}>0,\alpha>0,$ and $\beta>0$ with $\alpha\leq 1\leq\beta$ such that
    $c_{*}\left([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\right)$$\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad x\in\mathbb{R}^{n};$
  3. (iii)
    for any sequence of points $x_{k}$ in $\mathbb{R}^{n}$ one has
    1. (iii1)
      if $\psi(x_{k})\to 0,$ then $[f_{*}-f(x_{k})]_{+}\to 0;$
    2. (iii2)
      if $f(x_{k})\to-\infty,$ then $\psi(x_{k})\to\infty.$
When these equivalent properties hold, one has moreover for all $c>c_{*}$ that
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big{\{}f(x)+c\left([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\right)\big{\}}.$
Proof.
The equivalence between (i) and (i) as well as the last conclusion follow from Theorem 3.1, while the equivalence of (ii) and (iii) is known from Proposition 5.1. ∎
The term $[\psi(x)]^{\beta}$ cannot be omitted as shown in the following example.
Example 5.3.
Let $n:=2$ and consider the polynomial functions $f(x_{1},x_{2}):=-x_{1}^{2}-x_{2}^{2}$ and $\psi(x_{1},x_{2}):=x_{1}^{2}+x_{2}^{4}.$ By definition, we have that $\psi$ is a residual function for $S:=\{(0,0)\}$ and $f_{*}:=\inf_{(x_{1},x_{2})\in S}f(x_{1},x_{2})=0.$ Moreover, it is not hard to see that there are no constants $c>0$ and $\alpha>0$ such that
$c[\psi(x_{1},x_{2})]^{\alpha}$$\geq$ $[f_{*}-f(x_{1},x_{2})]_{+}\quad\textrm{ for all }\quad(x_{1},x_{2})\in\mathbb{R}^{2}.$
On the other hand, it holds that
$[\psi(x_{1},x_{2})]^{\frac{1}{2}}+\psi(x_{1},x_{2})$$\geq$ $[f_{*}-f(x_{1},x_{2})]_{+}\quad\textrm{ for all }\quad(x_{1},x_{2})\in\mathbb{R}^{2},$
which, together with Theorem 5.3, implies that for all $c>1,$
$\inf_{(x_{1},x_{2})\in S}f(x_{1},x_{2})$ $=$ $\inf_{(x_{1},x_{2})\in\mathbb{R}^{2}}\left\{f(x_{1},x_{2})+c\left([\psi(x_{1},x_{2})]^{\frac{1}{2}}+\psi(x_{1},x_{2})\right)\right\}.$
Remark 5.3.
(i) It is worth noting here that the exponents $\alpha$ and $\beta$ can take, respectively, to be $\frac{1}{N}$ and $N$ for some integer $N>0.$
(ii) To incorporate an abstract constraint of the form $x\in X$ with $X$ being a semi-algebraic subset of $\mathbb{R}^{n},$ we simply replace $f$ by $f+\delta_{X}(\cdot).$
In the rest of this subsection, we consider the problem (P) with $f:=f_{0}+\delta_{X},$ $f_{0}\colon X\to\mathbb{R}$ Lipschitz, $X\subset\mathbb{R}^{n}$ and $S\subset X$ nonempty and semi-algebraic. Let $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ be a residual function for $S$ which is semi-algebraic.
Theorem 5.4.
Let the following two conditions be satisfied
  1. (C1)
    if $\psi(x_{k})\to 0$ with $x_{k}\in X,$ then $\mathrm{dist}(x_{k},S)\to 0;$
  2. (C2)
    the distance function $\mathrm{dist}(\cdot,S)$ is bounded from above on $X.$
Then there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that for all $c>c_{*},$
$\inf_{x\in S}f_{0}(x)$ $=$ $\inf_{x\in X}\big{\{}f_{0}(x)+c\,[\psi(x)]^{\alpha}\big{\}},$
$\mathrm{argmin}_{S}f_{0}(x)$ $=$ $\mathrm{argmin}_{X}\big{\{}f_{0}(x)+c\,[\psi(x)]^{\alpha}\big{\}}.$
Proof.
By Theorem 3.1, it suffices to show the first equality.
According to Corollary 5.1, there exist some constants $c_{1}>0$ and $\alpha\in(0,1]$ such that
$c_{1}\,[\psi(x)]^{\alpha}$$\geq$ $\mathrm{dist}(x,S)\quad\textrm{ for all }\quad x\in X.$
Let $L>0$ be the Lipschitzian modulus of $f_{0}$ on X; i.e., for all $x,x^{\prime}\in X,$
$|f_{0}(x)-f_{0}(x^{\prime})|$$\leq$ $L\|x-x^{\prime}\|.$
Also let $c_{*}$ be an arbitrary scalar greater than $c_{1}L.$ Let $c>c_{*}$ and $x\in X$ an arbitrary point. Take any $\epsilon>0.$ By definition, there exists $x^{\prime}\in S$ such that
$\|x-x^{\prime}\|$$\leq$ $\mathrm{dist}(x,S)+\epsilon.$
On the other hand, we have $f_{0}(x^{\prime})\geq\inf_{S}f_{0}$ and $f_{0}(x)-f_{0}(x^{\prime})\geq-L\|x-x^{\prime}\|.$ Thus,
$f_{0}(x)+c\,[\psi(x)]^{\alpha}$ $=$ $f_{0}(x^{\prime})+\big{(}f_{0}(x)-f_{0}(x^{\prime})\big{)}+c\,[\psi(x)]^{\alpha}$
$\geq$ $f_{0}(x^{\prime})-L\|x-x^{\prime}\|+\frac{c}{c_{1}}\,\mathrm{dist}(x,S)$
$\geq$ $f_{0}(x^{\prime})-L\|x-x^{\prime}\|+\frac{c}{c_{1}}\big{(}\|x-x^{\prime}\|-\epsilon\big{)}$
$=$ $f_{0}(x^{\prime})+\left(\frac{c}{c_{1}}-L\right)\|x-x^{\prime}\|-\frac{c}{c_{1}}\epsilon$
$\geq$ $\inf_{S}f_{0}-\frac{c}{c_{1}}\epsilon.$
Taking the limit as $\epsilon\searrow 0$ yields the inequality
$f_{0}(x)+c\,[\psi(x)]^{\alpha}$$\geq$ $\inf_{S}f_{0}.$
Since $x$ was arbitrary in $X,$ we get
$\inf_{x\in S}f_{0}(x)$$\geq$ $\inf_{x\in X}\left\{f_{0}(x)+c\,[\psi(x)]^{\alpha}\right\}.$
The opposite inequality holds trivially, so the desired equality follows. ∎
Remark 5.4.
(i) A semi-algebraic version of [27, Theorem 2.1.2] is derived directly from Theorem 5.4 because the conditions of the theorem hold automatically when $X$ is compact and $\psi$ is continuous.
(ii) Assume that $S$ is given by
$S$ ::=:= $\{x\in\mathbb{R}^{n}\mid g_{1}(x)\leq 0,\ldots,g_{m}(x)\leq 0\},$
where $(g_{1},\ldots,g_{m})\colon\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a convenient and non-degenerate polynomial mapping (see [19, Section 3.7]. Certainly the semi-algebraic function
$\psi\colon\mathbb{R}^{n}\to\mathbb{R},\quad x\mapsto\max\{g_{1}(x),\ldots,g_{m}(x),0\},$
is residual of $S.$ Moreover, in view of [19, Theorem 3.7], there exist constants $c_{*}>0$ and $\alpha\in(0,1)$ such that
$c_{*}\,\big{(}[\psi(x)]^{\alpha}+[\psi(x)]\big{)}$$\geq$ $\mathrm{dist}(x,S)\quad\textrm{ for all }\quad x\in\mathbb{R}^{n}.$
Then by the same argument used earlier, we can see that for all $c>c_{*},$
$\inf_{x\in S}f_{0}(x)$ $=$ $\inf_{x\in X}\left\{f_{0}(x)+c\big{(}[\psi(x)]^{\alpha}+[\psi(x)]\right)\big{\}},$
$\mathrm{argmin}_{S}f_{0}(x)$ $=$ $\mathrm{argmin}_{X}\big{\{}f_{0}(x)+c\left([\psi(x)]^{\alpha}+[\psi(x)]\right)\big{\}}.$
In the general case, we have the following result.
Theorem 5.5.
Assume that for any sequence of points $x_{k}$ in $X$ the following properties hold
  1. (C1)
    if $\psi(x_{k})\to 0,$ then $\mathrm{dist}(x_{k},S)\to 0;$
  2. (C2)
    if $\mathrm{dist}(x_{k},S)\to\infty,$ then $\psi(x_{k})\to\infty.$
Then there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that for all $c>c_{*},$
$\inf_{x\in S}f_{0}(x)$ $=$ $\inf_{x\in X}\left\{f_{0}(x)+c\big{(}[\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\right)\big{\}},$
$\mathrm{argmin}_{S}f_{0}(x)$ $=$ $\mathrm{argmin}_{X}\big{\{}f_{0}(x)+c\left([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\right)\big{\}}.$
Proof.
The proof is similar to the one of Theorem 5.4 with taking into account Proposition 5.1 instead of Corollary 5.1. ∎
6. Conclusion
In this paper, we have presented necessary and sufficient conditions for the exactness of penalty functions in optimization problems whose constraint sets are not necessarily bounded. The conditions are given in terms of properties of the objective and residual functions of the problems in question. Our results generalize and improve some known facts in the literature on exact penalty functions. It would be nice to have applications of the results given here in optimization. This will be studied in the future research.
Acknowledgments
A part of this work was done while the second author and the third author were visiting Academy for Advanced Interdisciplinary Studies, Northeast Normal University, Changchun, China; they are grateful to the Academy for its hospitality and support.