### v2

# Exact penalty functions in optimization with unbounded constraint sets

Abstract.
This paper identifies necessary and sufficient conditions for the exactness of penalty functions in optimization problems whose constraint sets are not necessarily bounded. The case where the data of problems is locally Lipschitz, semi-algebraic or non-degenerate polynomials is studied in detail. The conditions are given in terms of properties of the objective and residual functions of the problems in question. The obtained results generalize and improve some known results in the literature on exact penalty functions.
Key words and phrases:
constrained optimization, exact penalty, calmness, semi-algebraic, Łojasiewicz inequality, non-degeneracy, Newton polyhedron
2010 Mathematics Subject Classification:
Primary 90C26; Secondary 49J52, 14P10
The work was partially supported by the Chinese National Natural Science Foundation under grant numbers 12471478, 12371300.
Corresponding author
1. Introduction
Constrained optimization problems arise in various fields, including engineering, economics, and machine learning, where the goal is to optimize an objective function subject to a set of constraints. Problems with complicated constraints are very difficult to deal with. A common approach for solving such problems is the usage of penalty function methods, wherein a constrained optimization problem is transformed into a sequence of unconstrained optimization problems whose solutions ideally converge to a solution of the original constrained problem; the unconstrained problems are formed by bringing the constraints into the objective function via a residual function and a scalar penalty parameter. Among these methods, exact penalty functions are particularly important because they allow us to solve a constrained optimization problem by solving a single unconstrained optimization problem. More precisely, a penalty function is said to be exact (or to have the exact penalty property) if there is a penalty parameter for which the corresponding unconstrained optimization problem and the original constrained optimization problem have the same optimal value and the same optimal solutions.
The concept of exact penalization traces back to the seminal works of Eremin [20] and Zangwill [46] (see also [40]). There is a huge literature on all aspects of the theory and applications of exact penalty functions; for more details, we refer the reader to the monographs [11, 21, 27, 28, 34, 37, 43, 48], the surveys [6, 8, 22] and the recent papers [23, 29, 41, 45] with the references therein.
In this paper we would like to identify necessary and sufficient conditions for a penalty function to be exact; the conditions are given in terms of properties of the objective and residual functions of the problems in question. As far as we know, for optimization problems whose constraint sets are unbounded, there are not many such studies; see [3, 10, 18, 47, 48, 49].
Contributions
We mainly study the exactness of penalty functions for constrained optimization problems with potentially unbounded constraint sets. Among other things, our main contributions are given in several steps:
  • For problems defined by locally Lipschitz functions, under some regularity conditions, a new necessary and sufficient condition for a penalty function to be exact is derived (see Theorem 4.1); this result improves [47, Theorems 1.5 and 3.1] as well as [49, Theorems 1.3] (in the finite-dimensional setting).
  • Based on the study of Łojasiewicz inequalities on unbounded sets, exact penalty functions for constrained optimization problems with semi-algebraic data are proposed. In the semi-algebraic setting, the obtained results generalize [44, Theorem 1], [12, Theorem 3.1] and [34, Theorem 2.1.2], wherein optimization problems over compact sets with subanalytic data are studied. It should be noted that the cited theorems could fail to hold when the compact assumption is absent (see Examples 5.3 and 5.5).
  • Based on the theory of Newton polyhedra, exact penalty functions for non-degenerate polynomial optimization problems are provided (see Theorems 6.1, 6.2 and 6.3). These results, together with ones in [15, 17, 24, 25, 38], suggest that the class of polynomial mappings, which is non-degenerate, may offer an appropriate domain on which the machinery of polynomial optimization works with full efficiency.
At this point, we would like to note the following facts:
  • the closedness of constraint sets and the continuity of objective functions, except for Theorems 4.1 and 6.1, are not required;
  • the boundedness of constraint sets is not imposed;
  • the existence of optimal solutions is not assumed.
We confine our study to the finite-dimensional case for two reasons. First of all, to lighten the exposition, we would like to concentrate on the basic ideas without technical and notational complications. Furthermore, the results obtained in Sections 5 and 6 are peculiar to finite-dimensions.
The tools used in this study come from variational analysis and semi-algebraic geometry.
The rest of the paper is organized as follows. Some definitions and preliminary results from variational analysis and semi-algebraic geometry are recalled in Section 2. The results and their proofs are presented in Sections 3, 4, 5 and 6. Conclusions are given in Section 7.
2. Preliminaries
2.1. Notation
Throughout this work we deal with the Euclidean space $\mathbb{R}^{n}$ equipped with the usual scalar product $\langle\cdot,\cdot\rangle$ and the corresponding norm $\|\cdot\|.$ We denote by $\mathbb{B}_{r}(x)$ the closed ball centered at $x$ with radius $r;$ when x{x} is the origin of $\mathbb{R}^{n}$ we write $\mathbb{B}_{r}$ instead of $\mathbb{B}_{r}({x}),$ and when $r=1$ we write $\mathbb{B}$ instead of $\mathbb{B}_{1}.$ We will adopt the convention that $\inf\emptyset=\infty$ and $\sup\emptyset=-\infty;$ the notation $x\to\infty$ means that $\|x\|\to\infty$.
Let $\mathbb{R}_{+}:=[0,\infty),\mathbb{R}_{-}:=(-\infty,0]$ and $\overline{\mathbb{R}}:=\mathbb{R}\cup\{\infty\}.$ For a real number $r,$ we write $r]_{+}:=\max\{r,0\}.$
For a nonempty set $\Omega\subset\mathbb{R}^{n},$ the closure, convex hull and conic hull of $\Omega$ are denoted, respectively, by $\mathrm{cl}\,{\Omega},$ $\mathrm{co}\,\Omega$ and $\mathrm{cone}\,\Omega.$ We will associate with $\Omega$ the distance function
$\mathrm{dist}(\cdot,\Omega)\colon\mathbb{R}^{n}\to\mathbb{R},\quad x\mapsto\inf_{x^{\prime}\in\Omega}\|x-x^{\prime}\|,$
and define the Euclidean projector of $x\in\mathbb{R}^{n}$ to $\Omega$ by
$\Pi_{\Omega}(x)$ $:=$ $\{x^{\prime}\in\Omega\ |\ \|x-x^{\prime}\|=\mathrm{dist}(x,\Omega)\}.$
The indicator function $\delta_{\Omega}\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$ of $\Omega$ is defined by
$\delta_{\Omega}(x)$ $:=$ $\begin{cases}0&\textrm{ if }x\in\Omega,\\ \infty&\textrm{ otherwise.}\end{cases}$
For an extended real-valued function $f\colon\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}},$ we denote its effective domain, graph, and epigraph by, respectively,
$\mathrm{dom}f$ $:=$ $\{x\in\mathbb{R}^{n}\ |\ f(x)<\infty\},$
$\mathrm{gph}f$ $:=$ $\{(x,y)\in\mathbb{R}^{n}\times\mathbb{R}\ |\ f(x)=y\},$
$\mathrm{epi}f$ $:=$ $\{(x,y)\in\mathbb{R}^{n}\times\mathbb{R}\ |\ f(x)\leq y\}.$
We call $f$ a proper function if $f(x)<\infty$ for at least one $x\in\mathbb{R}^{n},$ or in other words, if $\mathrm{dom}f$ is a nonempty set. The function $f$ is said to be lower semicontinuous if its epigraph is a closed set.
For an extended real-valued function $f\colon\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ and a set $\Omega\subset\mathbb{R}^{n},$ we let
$\inf_{\Omega}f$ $:=$ $\inf_{x\in\Omega}f(x)\ :=\ \inf\{f(x)\mid x\in\Omega\}.$
We introduce notation also for the set of points $x$ where the minimum of $f$ over $\Omega$ is regarded as being attained:
$\mathrm{argmin}_{\Omega}f(x)$ $:=$ $\begin{cases}\{x\in\Omega\mid f(x)=\inf_{x\in\Omega}f(x)\}&\textrm{ if }\inf_{x\in\Omega}f(x)\neq\infty,\\ \emptyset&\textrm{ otherwise.}\end{cases}$
For a set-valued mapping $F\colon\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m},$ we denote its graph by
$\mathrm{gph}F$ $:=$ $\{(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{m}\mid y\in F(x)\};$
the Painlevé–Kuratowski outer limit of $F$ at $x\in\mathbb{R}^{n}$ is defined by
$\mathop{{\rm Lim}\,{\rm sup}}_{x^{\prime}\to{x}}F(x)$ $:=$ $\{y\in\mathbb{R}^{m}\mid\exists x_{k}\to{x},\exists y_{k}\in F(x_{k}),y_{k}\to y\}.$
2.2. Normal cones, subdifferentials and coderivatives
Here we recall some definitions and results of variational analysis, which can be found in [36, 37, 43].
Definition 2.1.
Consider a set $\Omega\subset\mathbb{R}^{n}$ and a point ${x}\in\Omega.$ The limiting normal cone (also known as the basic or Mordukhovich normal cone) to $\Omega$ at x{x} is defined by
$N_{\Omega}(x)$ $:=$ $\mathop{{\rm Lim}\,{\rm sup}}_{x^{\prime}\rightarrow x}\Big[\mathrm{cone}\big(x^{\prime}-\Pi_{\Omega}(x^{\prime})\big)\Big].$
If $x\not\in\Omega,$ we put ${N}_{\Omega}({x}):=\emptyset.$
Remark 2.1.
(i) It is well known that $N_{\Omega}({x})$ is a closed (possibly non-convex) cone.
(ii) If $\Omega$ is a manifold of class $C^{1},$ then the normal cone $N_{\Omega}({x})$ is equal to the normal space to $\Omega$ at x{x} in the sense of differential geometry; see [43, Example 6.8].
Definition 2.2.
Consider an extended real-valued function $f\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$ and a point ${x}\in\mathrm{dom}f.$ The limiting (or Mordukhovich) subdifferential of $f$ at x{x} is defined by
$\partial f({x})$ $:=$ $\{u\in\mathbb{R}^{n}\mid(u,-1)\in{N}_{\mathrm{epi}f}({x},f({x}))\}.$
If $x\not\in\mathrm{dom}f,$ we put $\partial f({x}):=\emptyset.$
Remark 2.2.
In [36, 37, 43] the reader can find equivalent analytic descriptions of the limiting subdifferential $\partial f({x})$ and comprehensive studies of it and related constructions. For convex $f,$ this subdifferential coincides with the convex subdifferential. Furthermore, if the function $f$ is of class $C^{1},$ then $\partial f({x})=\{\nabla f({x})\}.$
Definition 2.3.
Consider a set-valued mapping $F\colon\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ and a point $({x},{y})\in\mathrm{gph}F.$ The (basic) coderivative of $F$ at (x,y)({x},{y}) is the set-valued mapping $D^{*}F({x},{y})\colon\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{n}$ defined by
$D^{*}F({x},{y})(v)$ $:=$ $\{u\in\mathbb{R}^{n}\mid(u,-v)\in N_{\mathrm{gph}F}({x},{y})\}\quad\textrm{ for all }\quad v\in\mathbb{R}^{m}.$
The following facts are well known (see, for example, [37]).
Lemma 2.1.
Consider a nonempty closed set $\Omega\subset\mathbb{R}^{n}.$ We have for all ${x}\in\Omega,$
$N_{\Omega}({x})$ $=$ $\mathop{{\rm Lim}\,{\rm sup}}_{x^{\prime}\xrightarrow{\Omega}{x}}{N}_{\Omega}(x^{\prime}),$
where $x^{\prime}\xrightarrow{\Omega}{x}$ means that $x^{\prime}\rightarrow{x}$ with $x^{\prime}\in\Omega.$
Lemma 2.2.
Let $\Omega\subset\mathbb{R}^{n}$ be a nonempty closed set. Then
$\partial\mathrm{dist}(\cdot,\Omega)(x)$ $=$ $\begin{cases}N_{\Omega}(x)\cap\mathbb{B}&\quad\textrm{ if }x\in\Omega,\\ \frac{x-\Pi_{\Omega}(x)}{\mathrm{dist}(x,\Omega)}&\quad\textrm{ otherwise.}\end{cases}$
Lemma 2.3.
For a lower semicontinuous function $f\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$ and a point $\overline{x}\in\mathrm{dom}f,$ we have
$\partial f(\overline{x})$ $=$ $\mathop{{\rm Lim}\,{\rm sup}}_{x\xrightarrow{f}\overline{x}}{\partial}f(x),$
where $x\xrightarrow{f}\overline{x}$ means that $x\to\overline{x}$ and $f(x)\to f(\overline{x}).$
Lemma 2.4.
Let $f_{i}\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$, $i=1,\dots,m$ with $m\geq 2$, be lower semicontinuous at $\overline{x}\in\mathbb{R}^{n}$ and let all but one of these functions be Lipschitz continuous around $\overline{x}.$ Then the following inclusions hold
$\partial\left(f_{1}+\cdots+f_{m}\right)(\overline{x})$ $\subset$ $\partial f_{1}(\overline{x})+\cdots+\partial f_{m}(\overline{x}),$
$\partial(\max f_{i})(\overline{x})$ $\subset$ $\bigcup\mathrm{co}(\{v_{i}\mid i\in I(\overline{x})\}),$
where $I(\overline{x}):=\{i\in\{1,\ldots,m\}\mid f_{i}(\overline{x})=\max_{j}f_{j}(\overline{x})\}$ and the union is taken over all vectors $v_{i}\in\partial f_{i}(\overline{x})$ for $i\in I(\overline{x}).$
Lemma 2.5 (Fermat rule).
If a proper function $f\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$ has a local minimum at $\overline{x}\in\mathrm{dom}f,$ then $0\in\partial f(\overline{x}).$
Lemma 2.6.
Let $f\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be Lipschitz continuous around $\overline{x}\in\mathbb{R}^{n}$ with modulus $L.$ Then $\partial f(\overline{x})\neq\emptyset$ and
$\|v\|$ $\leq$ $L\quad\textrm{ for all }\quad v\in\partial f(\overline{x}).$
Lemma 2.7 (chain rule).
Let $g\colon\mathbb{R}^{n}\to{\mathbb{R}}^{m}$ be Lipschitz continuous around $\overline{x}\in\mathbb{R}^{n}$ and $h\colon\mathbb{R}^{m}\to{\mathbb{R}}$ be Lipschitz continuous around $\overline{y}:=g(\overline{x})\in\mathbb{R}^{m}.$ Then
$\partial(h\circ g)(\overline{x})$ $\subset$ $\bigcup_{w\in\partial h(\overline{y})}\partial\langle w,g\rangle(\overline{x}).$
Finally, the next lemma expresses the coderivative of a single-valued Lipschitz mapping $g\colon\mathbb{R}^{n}\to{\mathbb{R}}^{m}$ via the subdifferential of the scalarization $\langle w,g\rangle\colon\mathbb{R}^{n}\to{\mathbb{R}},x\mapsto\langle w,g(x)\rangle.$
Lemma 2.8.
Let $g\colon\mathbb{R}^{n}\to{\mathbb{R}}^{m}$ be Lipschitz continuous around $\overline{x}\in\mathbb{R}^{n}.$ Then
$D^{*}g(\overline{x})(w)$ $=$ $\partial\langle w,g\rangle(\overline{x})\quad\textrm{ for all }\quad w\in\mathbb{R}^{m}.$
2.3. Semi-algebraic geometry
Now, we recall some notions and results of semi-algebraic geometry, which can be found in [1, 5, 25].
Definition 2.4.
A subset $S$ of $\mathbb{R}^{n}$ is semi-algebraic if it is a finite union of sets of the form
$\{x\in\mathbb{R}^{n}\ |\ f_{i}(x)=0,\ i=1,\ldots,k;f_{j}(x)>0,\ j=k+1,\ldots,p\},$
where all $f_{i}$ are polynomials. In other words, $S$ is a union of finitely many sets, each defined by finitely many polynomial equalities and inequalities.
A function $f\colon S\rightarrow{\mathbb{R}\cup\{\infty\}}$ is said to be semi-algebraic if its graph
$\{(x,y)\in S\times\mathbb{R}\ |\ y=f(x)\}$
is a semi-algebraic set.
Example 2.1.
Each polyhedral set is the intersection of a finite number of half-spaces, and so is semi-algebraic.
A major fact concerning the class of semi-algebraic sets is its stability under linear projections.
Theorem 2.1 (Tarski–Seidenberg theorem).
The image of any semi-algebraic set $S\subset\mathbb{R}^{n}$ under a projection to any linear subspace of $\mathbb{R}^{n}$ is a semi-algebraic set.
Remark 2.3.
As an immediate consequence of the Tarski–Seidenberg theorem, we get the semi-algebraicity of any set $\{x\in A\ |\ \exists y\in B,(x,y)\in C\},$ provided that $A,B,$ and $C$ are semi-algebraic sets in the corresponding spaces. Also, $\{x\in A\ |\ \forall y\in B,(x,y)\in C\}$ is a semi-algebraic set as its complement is the union of the complement of $A$ and the set $\{x\in A\ |\ \exists y\in B,(x,y)\not\in C\}.$ Thus, if we have a finite collection of semi-algebraic sets, then any set obtained from them with the help of a finite chain of quantifiers on variables is also semi-algebraic.
The following well known lemmas will be of great importance for us (see, for example, [1, 5, 25]).
Lemma 2.9.
Every semi-algebraic set in $\mathbb{R}$ is a finite union of points and intervals.
Lemma 2.10 (curve selection lemma at infinity).
[…]
Lemma 2.11 (growth dichotomy lemma).
  1. (i)
    Let $f\colon(0,\epsilon)\rightarrow{\mathbb{R}}$ be a semi-algebraic function with $f(t)\neq 0$ for all $t\in(0,\epsilon).$ Then there exist constants $a\neq 0$ and $\alpha\in{\mathbb{Q}}$ such that $f(t)=at^{\alpha}+o(t^{\alpha})$ as $t\to 0^{+}.$
  2. (ii)
    Let $f\colon(r,+\infty)\rightarrow{\mathbb{R}}$ be a semi-algebraic function with $f(t)\neq 0$ for all $t\in(r,+\infty).$ Then there exist constants $b\neq 0$ and $\beta\in{\mathbb{Q}}$ such that $f(t)=bt^{\beta}+o(t^{\beta})$ as $t\to+\infty.$
Lemma 2.12 (monotonicity lemma).
Let $f\colon(a,b)\to\mathbb{R}$ be a semi-algebraic function. Then there are $a=a_{0}<a_{1}<\cdots<a_{s}<a_{s+1}=b$ such that, for each $i=0,\ldots,s,$ the restriction $f|_{(a_{i},a_{i+1})}$ is analytic and either constant, strictly increasing, or strictly decreasing.
3. Exact penalty functions: the general case
Throughout this paper, we consider the following constrained optimization problem
$\textrm{minimize }f(x)\quad\textrm{ subject to }x\in S,$ (P)
where $S$ is a nonempty subset of $\mathbb{R}^{n}$ and $f\colon\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is an extended real-valued function. To avoid triviality, we assume in what follows that the optimal value $f_{*}:=\inf_{x\in S}f(x)$ of the problem (P) is finite. In this section, we will present some global necessary and sufficient condition for a penalty function to be exact. Let us start with the following.
Definition 3.1.
A residual function of $S$ is a nonnegative valued function $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ such that $\psi(x)=0$ if and only if $x\in S.$
Remark 3.1.
If $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ is a residual function for $S,$ then for any function $\phi\colon\mathbb{R}^{n}\to(0,\infty)$ and any constants $\alpha>0$ and $\beta>0,$ one has the functions $\phi\cdot\psi,$ $\psi^{\alpha}$ and $\psi^{\alpha}+\psi^{\beta}$ are residual for $S.$
The following simple fact will play an important role in establishing exact penalty properties in optimization problems with unbounded constraint sets.
Theorem 3.1.
For any residual function $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ of $S,$ the following three properties are equivalent:
  1. (i)
    for all $c$ sufficiently large, one has
    $\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c\,\psi(x)\big\};$
  2. (ii)
    there exists a constant $c_{*}>0$ such that
    $\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c_{*}\,\psi(x)\big\};$
  3. (iii)
    there exists a constant $c_{*}>0$ such that
    $c_{*}\psi(x)$ $\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad x\in\mathbb{R}^{n}.$ (1)
When these equivalent properties hold, one has for all $c>c_{*}$ that
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big\{f(x)+c\,\psi(x)\big\}.$
Proof.
The implications (i) $\Rightarrow$ (ii) $\Rightarrow$ (iii) is obvious.
(iii) $\Rightarrow$ (i) Fix $c>c_{*}$ and take any $x\in\mathbb{R}^{n}.$ If $f(x)\geq f_{*},$ then $f(x)+c\,\psi(x)\geq f_{*};$ otherwise (1) gives
$f(x)+c\,\psi(x)$ $\geq$ $f(x)+c_{*}\,\psi(x)\ \geq\ f(x)+[f_{*}-f(x)]_{+}\ =\ f_{*}.$
In both cases we have
$f(x)+c\,\psi(x)$ $\geq$ $f_{*}.$
Since $x\in\mathbb{R}^{n}$ was chosen arbitrarily, we conclude that
$\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c\,\psi(x)\big\}$ $\geq$ $f_{*}.$
On the other hand, by definition, we have
$\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c\,\psi(x)\big\}$ $\leq$ $\inf_{x\in S}\big\{f(x)+c\,\psi(x)\big\}\ =\ \inf_{x\in S}f(x)\ =\ f_{*}.$
Therefore, (i) holds.
To show the last conclusion, assume that $x_{0}$ minimizes $f$ on $S.$ Then
$f(x_{0})+c\,\psi(x_{0})$ $=$ $f(x_{0})\ =\ f_{*},$
which, together with (i), yields that $x_{0}$ minimizes $f+c\psi$ on $\mathbb{R}^{n}.$
Conversely, assume that $x_{0}$ minimizes $f+c\psi$ on $\mathbb{R}^{n}.$ Since $c>c_{*}$ and the function $\psi$ is nonnegative, it holds that
$f(x_{0})+c_{*}\psi(x_{0})\leq f(x_{0})+c\,\psi(x_{0})$ $\leq$ $f_{*}.$
On the other hand, it follows from (1) that
$f(x_{0})+c_{*}\psi(x_{0})$ $\geq$ $f_{*}.$
Therefore,
$f(x_{0})+c_{*}\psi(x_{0})$ $=$ $f(x_{0})+c\,\psi(x_{0})\ =\ f_{*}.$
Since $c_{*}<c,$ then $\psi(x_{0})=0.$ Hence $x_{0}\in S$ and $f(x_{0})=f_{*}.$ Thus, $x_{0}$ minimizes $f$ on $S.$
Example 3.1.
Let $n:=1,S:=\{x\in\mathbb{R}\mid x<1\}$ and $f(x):=|x|+\delta_{\{x\neq 0\}}(x).$ By definition, the function $\psi\colon\mathbb{R}\to\mathbb{R}$ defined by
$\psi(x):=\begin{cases}\max\{x-1,0\}&\textrm{ if }x\neq 1,\\ 1&\textrm{ otherwise}\end{cases}$
is a residual function for $S.$ Fix any $c_{*}>0.$ It is easy to check that $f_{*}=0$ and
$c_{*}\psi(x)$ $\geq$ $0\ =\ [f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad x\in\mathbb{R}.$
By Theorem 3.1, for any $c>c_{*}$ we have
$\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}}\big\{f(x)+c\,\psi(x)\big\},$
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}}\big\{f(x)+c\,\psi(x)\big\}\ =\ \emptyset.$
Also observe that the constraint set $S$ is neither closed nor bounded and the objective function $f$ is not lower semicontinuous.
The results presented in the rest of this section are inspired by the works of Clarke [9, Section 6.4] and Burke [7, 8]. To begin with, assume that
$S$ $:=$ $\{x\in\mathbb{R}^{n}\mid g(x)\in C\},$
where $g\colon\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is a mapping and $C$ is a nonempty subset of $\mathbb{R}^{m}.$ Then the problem (P) can be rewritten as
$\textrm{minimize }f(x)\quad\textrm{ subject to }g(x)\in C.$ (P0)
In this case, an equivalent relation between the exact penalization and the global calmness can be stated as follows.
Theorem 3.2.
The following two properties are equivalent:
  1. (i)
    for all $c$ sufficiently large one has
    $\inf_{g(x)\in{C}}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c\,\mathrm{dist}\big(g(x),{C}\big)\big\};$
  2. (ii)
    the problem (P0) is globally calm in the sense that there exists a constant $c_{*}>0$ such that for every pair $(\overline{x},\overline{u})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ with $g(\overline{x})\in{C}+\overline{u},$ one has
    $\inf_{g(x)\in{C}}f(x)$ $\leq$ $f(\overline{x})+c_{*}\,\|\overline{u}\|.$
When these equivalent properties hold, one has for all $c>c_{*}$ that
$\mathrm{argmin}_{g(x)\in C}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big\{f(x)+c\,\mathrm{dist}(g(x),C)\big\}.$
Proof.
It suffices to show the equivalence of (i) with (ii) since the last conclusion is a direct consequence of Theorem 3.1.
(i) $\Rightarrow$ (ii) Assume that
$\inf_{g(x)\in{C}}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c_{*}\,\mathrm{dist}\big(g(x),{C}\big)\big\}$
for some $c_{*}>0.$ Let $(\overline{x},\overline{u})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ be such that $g(\overline{x})\in{C}+\overline{u}.$ Then
$\inf_{g(x)\in{C}}f(x)$ $\leq$ $f(\overline{x})+c_{*}\,\mathrm{dist}\big(g(\overline{x}),{C}\big)$
$=$ $f(\overline{x})+c_{*}\,\inf\big\{\|g(\overline{x})-y\|\mid y\in{C}\big\}$
$\leq$ $f(\overline{x})+c_{*}\,\inf\big\{\|g(\overline{x})-(y+{\overline{u}})\|+\|\overline{u}\|\mid y\in{C}\big\}$
$=$ $f(\overline{x})+c_{*}\,\mathrm{dist}\big(g(\overline{x}),{C}+{\overline{u}}\big)+c_{*}\,\|\overline{u}\|$
$=$ $f(\overline{x})+c_{*}\,\|\overline{u}\|,$
which gives the desired result.
(ii) $\Rightarrow$ (i) Let $\overline{x}\in\mathbb{R}^{n}$ and take any $\epsilon>0.$ There exists a point $y\in{C}$ such that
$\|g(\overline{x})-y\|$ $\leq$ $\mathrm{dist}\big(g(\overline{x}),{C}\big)+\epsilon.$
Let $\overline{u}:=g(\overline{x})-y.$ Then $g(\overline{x})=y+\overline{u}\in{C}+\overline{u}.$ By hypothesis, we get
$\inf_{g(x)\in{C}}f(x)$ $\leq$ $f(\overline{x})+c_{*}\,\|\overline{u}\|$
$\leq$ $f(\overline{x})+c_{*}\,\mathrm{dist}\big(g(\overline{x}),{C}\big)+c_{*}\epsilon.$
Taking the limit as $\epsilon\searrow 0$ yields the inequality
$\inf_{g(x)\in{C}}f(x)$ $\leq$ $f(\overline{x})+c_{*}\,\mathrm{dist}\big(g(\overline{x}),{C}\big).$
Since $\overline{x}$ was arbitrary in $\mathbb{R}^{n},$ it follows that
$\inf_{g(x)\in{C}}f(x)$ $\leq$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c_{*}\,\mathrm{dist}\big(g(x),{C}\big)\big\}.$
Since the opposite inequality holds trivially, the desired equality follows. ∎
To study the exactness of penalty functions, we can imbed the problem (P0) into a parametrized family of mathematical programs. For each $u\in\mathbb{R}^{m}$ consider the perturbed problem
$\textrm{minimize }f(x)\quad\textrm{ subject to }g(x)\in C+u.$
Accordingly, we define the optimal value function $V\colon\mathbb{R}^{m}\to\mathbb{R}\cup\{\pm\infty\}$ by
$V(u)$ $:=$ $\begin{cases}\inf_{g(x)\in C+u}f(x)&\textrm{ if }\{x\in\mathbb{R}^{n}\mid g(x)\in C+u\}\neq\emptyset,\\ \infty&\textrm{ otherwise.}\end{cases}$
Corollary 3.1.
If there exists a constant $c_{*}>0$ such that
$\inf_{g(x)\in{C}}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c_{*}\,\mathrm{dist}\big(g(x),{C}\big)\big\},$
then
$\liminf_{u\to 0}\frac{V(u)-V(0)}{\|u\|}$ $>$ $-\infty.$
The converse holds true if $f$ is bounded from below on $\mathbb{R}^{n}.$
Proof.
Assume that
$\inf_{g(x)\in{C}}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c_{*}\,\mathrm{dist}\big(g(x),{C}\big)\big\}.$
Take any $u\in\mathbb{R}^{m}.$ By Theorems 3.1 and 3.2, if there exists a point $\overline{x}\in\mathbb{R}^{n}$ with $g(\overline{x})\in{C}+u,$ then
$\inf_{g(x)\in{C}}f(x)$ $\leq$ $f(\overline{x})+c_{*}\,\|u\|,$
which easily yields that
$V(0)$ $\leq$ $V(u)+c_{*}\,\|u\|.$
This inequality is still valid when there is no point $\overline{x}$ with $g(\overline{x})\in{C}+u.$ Since $u\in\mathbb{R}^{m}$ was arbitrary, we get the desired conclusion
$\liminf_{u\to 0}\frac{V(u)-V(0)}{\|u\|}$ $\geq$ $-c_{*}\ >\ -\infty.$
Now assume that $f$ is bounded from below on $\mathbb{R}^{n}$ and the converse is not true. By Theorem 3.2, there exists a sequence $(x_{k},u_{k})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ with $g(x_{k})\in C+u_{k}$ such that
$\inf_{g(x)\in{C}}f(x)$ $>$ $f(x_{k})+k\,\|u_{k}\|\ \geq\ \inf_{x\in\mathbb{R}^{n}}f(x)+k\,\|u_{k}\|.$
Certainly $u_{k}\neq 0$ and $u_{k}\to 0$ as $k\to\infty.$ Moreover, we have
$0$ $>$ $\big[f(x_{k})-\inf_{g(x)\in{C}}f(x)\big]+k\,\|u_{k}\|\ \geq\ V(u_{k})-V(0)+k\,\|u_{k}\|.$
Hence
$0$ $>$ $\frac{V(u_{k})-V(0)}{\|u_{k}\|}+k,$
which is impossible for $k$ sufficiently large. ∎
The converse of the corollary above is not true if $f$ is not bounded from below on $\mathbb{R}^{n};$ see Example 5.2(ii).
Remark 3.2.
The results presented in this section can be applied to an abstract constraint of the form $x\in X$ with $X$ being a subset of $\mathbb{R}^{n}$ by substituting $f$ by $f+\delta_{X}(\cdot).$
4. Exact penalty functions: the locally Lipschitz case
This section is inspired by the work of Zaslavski [47, 49]. Indeed, consider the problem (P) with the objective function $f$ being locally Lipschitz continuous and the constraint set $S$ being given by
$S$ $:=$ $\{x\in\mathbb{R}^{n}\ |\ g(x)\in C\},$
where $g\colon\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a locally Lipschitz mapping and $C$ is a nonempty closed subset of $\mathbb{R}^{m}.$ Assume that the optimal value $f_{*}:=\inf_{g(x)\in C}f(x)$ is finite. Under regularity conditions, local and at infinity, we will give a necessary and sufficient condition for the following to hold
$\inf_{g(x)\in C}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c\,\mathrm{dist}(g(x),C)\big\}$
for all $c$ sufficiently large.
Recall that for each vector $w\in\mathbb{R}^{m},$ the function $\langle w,g\rangle\colon\mathbb{R}^{n}\to\mathbb{R}$ is defined by
$\langle w,g\rangle(x)$ $:=$ $\langle w,g(x)\rangle\quad\textrm{ for }\quad x\in\mathbb{R}^{n}.$
Let us start with the following.
Definition 4.1.
We say that the (generalized) Mangasarian–Fromovitz constraint qualification (MFCQ) holds at $x\in S$ if there is no unit vector $w\in N_{C}(g(x))$ such that $0\in\partial\langle w,g\rangle(x).$
Remark 4.1.
When $C=\{0\}\times(\mathbb{R}_{-})^{m-p}\subset\mathbb{R}^{m}$ and $g$ is of class $C^{1},$ the above definition reduces to the traditional one; see [35].
Let $K_{\infty}(f,S)$ be the set of all $t\in\mathbb{R}$ for which there exists a sequence $x_{k}\in\mathbb{R}^{n}\setminus S$ satisfying the following conditions:
$\lim_{k\to\infty}\|x_{k}\|=\infty,\quad\lim_{k\to\infty}\|x_{k}\|\nu(x_{k})=0,\quad\lim_{k\to\infty}f(x_{k})=t,\quad\textrm{ and }\quad\lim_{k\to\infty}\mathrm{dist}(g(x_{k}),C)=0.$
Here and in the following we put
$\nu(x)$ $:=$ $\inf\|\lambda u+(1-\lambda)v\|\quad\textrm{ for }\quad x\in\mathbb{R}^{n}\setminus S$
with the infimum being taken over all real numbers $\lambda\in(0,1)$ and vectors $u\in\partial f(x),v\in\partial\langle w,g\rangle(x),$ and $w\in\frac{g(x)-\Pi_{C}(g(x))}{\mathrm{dist}(g(x),C)}.$
Remark 4.2.
(i) The definition of $K_{\infty}(f,S)$ is strongly related to the weak Palais–Smale condition (see, for example, [31, 42]).
(ii) When $C=\{0\}\subset\mathbb{R}^{m}$ and $f$ and $g$ are of class $C^{1},$ we have that
$\nu(x)$ $=$ $\inf\big\{\|\lambda\nabla f(x)+(1-\lambda)\nabla\|g\|(x)\|\mid\lambda\in(0,1)\big\}.$
(iii) If for some $c_{*}>0$ we have
$-\infty\ <\ \inf_{g(x)\in C}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c_{*}\,\mathrm{dist}(g(x),C)\big\},$
then the function
$\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{\infty\},\quad x\mapsto f(x)+c_{*}\,\mathrm{dist}(g(x),C),$
must be bounded from below. Conversely, we have the following result, which improves [47, Theorems 1.5 and 3.1] and [49, Theorems 1.3] (in the finite-dimensional setting).
Theorem 4.1.
Let the following two assumptions hold
  1. (A1)
    the (MFCQ) is satisfied at every global minimum of $f$ on $S;$
  2. (A2)
    the inclusion $K_{\infty}(f,S)\subset(f_{*},\infty)$ is valid.
Then the following properties are equivalent:
  1. (i)
    for all $c$ sufficiently large, one has
    $\inf_{g(x)\in C}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c\,\mathrm{dist}(g(x),C)\big\};$
  2. (ii)
    there exists a constant $c_{*}>0$ such that
    $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c_{*}\mathrm{dist}(g(x),C)\big\}$ $>$ $-\infty.$
When these equivalent properties hold, one has for any $c>c_{*}$ that
$\mathrm{argmin}_{g(x)\in C}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big\{f(x)+c\,\mathrm{dist}(g(x),C)\big\};$
moreover, if $\overline{x}$ is a global minimizer for the problem (P0), then there exists a vector $w\in N_{C}(g(\overline{x}))\cap\mathbb{B}$ satisfying the following optimality condition
$0$ $\in$ $\partial f(\overline{x})+c\,\partial\langle w,g\rangle(\overline{x}).$
To prove the theorem, we need the following simple fact.
Lemma 4.1.
Let $x_{k},v_{k}\in\mathbb{R}^{n}$ and $w_{k}\in\mathbb{R}^{m}$ be sequences converging to $\overline{x},\overline{v}$ and $\overline{w},$ respectively, such that $v_{k}\in\partial\langle w_{k},g\rangle(x_{k})$ for all $k.$ Then $\overline{v}\in\partial\langle\overline{w},g\rangle(\overline{x}).$
Proof.
Since $v_{k}\in\partial\langle w_{k},g\rangle(x_{k}),$ it follows from Lemma 2.8 that $v_{k}\in D^{*}g(x_{k})(w_{k}),$ which by definition is equivalent to $(v_{k},-w_{k})\in N_{\mathrm{gph}g}(x_{k},g(x_{k})).$ Letting $k\to\infty$ and applying Lemma 2.1, we get $(\overline{v},-\overline{w})\in N_{\mathrm{gph}g}(\overline{x},g(\overline{x})),$ and so $\overline{v}\in D^{*}g(\overline{x})(\overline{w}).$ By Lemma 2.8 again, $\overline{v}\in\partial\langle\overline{w},g\rangle(\overline{x}),$ as required. ∎
Proof of Theorem 4.1.
We first prove the equivalence of (i) with (ii). To see this, define the function $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ by
$\psi(x)$ $:=$ $\mathrm{dist}(g(x),C)\quad\textrm{ for }\quad x\in\mathbb{R}^{n},$
which is locally Lipschitz continuous and residual for $S.$ By Theorem 3.1, it suffices to show the equivalence of (ii) with the following property:
  1. ($\star$)
    there exists a constant $c_{*}>0$ such that
    $c_{*}\psi(x)$ $\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad x\in\mathbb{R}^{n}.$
Indeed, the ($\star$) $\Rightarrow$ (ii) part is trivial. For the (ii) $\Rightarrow$ ($\star$) part we likewise proceed by contradiction. Let $c_{*}>0$ be such that
$\inf_{x\in\mathbb{R}^{n}}\left\{f(x)+c_{*}\psi(x)\right\}$ $>$ $-\infty,$
and suppose that for each $k>c_{*},$ there exists $x_{k}\in\mathbb{R}^{n}$ such that
$k\psi(x_{k})$ $<$ $[f_{*}-f(x_{k})]_{+}.$
Since the function $\psi$ is nonnegative, $f_{*}-f(x_{k})>0.$ Moreover, if $\psi(x_{k})=0$ then $x_{k}\in S$ and so $f_{*}\leq f(x_{k}),$ a contradiction. Therefore
$0$ $<$ $k\psi(x_{k})\ <\ f_{*}-f(x_{k}).$ (2)
Define the function $F_{k}\colon\mathbb{R}^{n}\to\mathbb{R},x\mapsto F_{k}(x),$ by
$F_{k}(x)$ $:=$ $f(x)+k\psi(x),$
which is locally Lipschitz continuous and satisfies
$-\infty\ <\ \mathfrak{m}:=\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c_{*}\psi(x)\big\}$ $\leq$ $\inf_{x\in\mathbb{R}^{n}}F_{k}(x)\ <\ f_{*},$
where the last inequality follows from (2).
Put
$\varepsilon_{k}:=f_{*}-\inf_{x\in\mathbb{R}^{n}}F_{k}(x)\quad\textrm{ and }\quad\lambda_{k}:=\tfrac{\|x_{k}\|+1}{2}.$
Then $0<\varepsilon_{k}<f_{*}-\mathfrak{m}$ and we deduce from (2) that
$F_{k}(x_{k})$ $<$ $f_{*}\ =\ \inf_{x\in\mathbb{R}^{n}}F_{k}(x)+\varepsilon_{k}.$
By the Ekeland variational principle (see [19]), there exists $x_{k}^{\prime}\in\mathbb{R}^{n}$ having the following conditions:
  1. (c1)
    $F_{k}(x_{k}^{\prime})\leq F_{k}(x_{k})<f_{*}=\inf_{x\in\mathbb{R}^{n}}F_{k}(x)+\varepsilon_{k},$
  2. (c2)
    $\|x_{k}^{\prime}-x_{k}\|\leq\lambda_{k},$ and
  3. (c3)
    $F_{k}(x_{k}^{\prime})\leq F_{k}(x)+\frac{\varepsilon_{k}}{\lambda_{k}}\|x-x_{k}^{\prime}\|$ for all $x\in\mathbb{R}^{n}$.
Note that $\psi(x_{k}^{\prime})>0$ because otherwise we would have $x_{k}^{\prime}\in S,$ so that
$F_{k}(x_{k}^{\prime})$ $=$ $f(x_{k}^{\prime})+k\psi(x_{k}^{\prime})\ =\ f(x_{k}^{\prime})\ \geq\ f_{*}$
in contradiction to the condition (c1). Moreover, we have
$\mathfrak{m}$ $\leq$ $f(x_{k}^{\prime})+{c_{*}}\psi(x_{k}^{\prime})\ \leq\ f(x_{k}^{\prime})+k\psi(x_{k}^{\prime})\ <\ f_{*}.$ (3)
It follows that
$\mathfrak{m}+(k-c_{*})\psi(x_{k}^{\prime})$ $\leq$ $f_{*}.$
Consequently, $\psi(x_{k}^{\prime})$ tends to 0 as $k\to\infty.$ This when combined with (3) implies that the sequence $f(x_{k}^{\prime})$ is bounded and has all its cluster points in the interval $\mathfrak{m},f_{*}].$
From the condition (c2), we easily deduce that
$\frac{\|x_{k}\|-1}{2}$ $\leq$ $\|x_{k}^{\prime}\|\ \leq\ \frac{3\|x_{k}\|+1}{2}.$ (4)
In particular, the sequence $x_{k}$ is bounded if and only if the sequence $x_{k}^{\prime}$ is bounded.
On the other hand, by the condition (c3), $x_{k}^{\prime}$ is a minimizer of the locally Lipschitz function
$\mathbb{R}^{n}\to\mathbb{R},\quad x\mapsto F_{k}(x)+\frac{\varepsilon_{k}}{\lambda_{k}}\|x-x_{k}^{\prime}\|.$
Lemma 2.5 (Fermat rule) when combined with Lemma 2.4 (sum rule) yields
$0$ $\in$ $\partial\left[F_{k}+\frac{\varepsilon_{k}}{\lambda_{k}}\|\cdot-x_{k}^{\prime}\|\right](x_{k}^{\prime})$
$\subset$ $\partial f(x_{k}^{\prime})+k\partial\psi(x_{k}^{\prime})+\frac{\varepsilon_{k}}{\lambda_{k}}\mathbb{B}.$
Note that $g(x_{k}^{\prime})\not\in C$ because $\psi(x_{k}^{\prime})>0.$ Moreover, we know from Lemmas 2.2 and 2.7 that
$\partial\psi(x_{k}^{\prime})$ $\subset$ $\bigcup_{w}\partial\langle w,g\rangle(x_{k}^{\prime}),$
where the union is taken on all vectors $w$ belonging to the set $\partial\psi(g(x_{k}^{\prime})),$ which is
$\frac{g(x_{k}^{\prime})-\Pi_{C}(g(x_{k}^{\prime}))}{\psi(x_{k}^{\prime})}.$
Therefore, we can find vectors $u_{k}\in\partial f(x_{k}^{\prime})$ and $v_{k}\in\partial\langle w_{k},g\rangle(x_{k}^{\prime})$ for some $w_{k}\in\partial\psi(g(x_{k}^{\prime}))$ such that
$0$ $\in$ $u_{k}+kv_{k}+\frac{\varepsilon_{k}}{\lambda_{k}}\mathbb{B}.$ (5)
There are two cases to be considered.
Case 1: the sequence $\{x_{k}\}$ is bounded.
By (4), the sequence $\{x_{k}^{\prime}\}$ is bounded too. Passing to a subsequence if necessary, we can suppose that $x_{k}^{\prime}$ converges to some point $\overline{x}\in\mathbb{R}^{n}.$ Certainly, $\psi(\overline{x})=0$ (equivalently, $\overline{x}\in S$) and $f(\overline{x})=f_{*},$ which imply that $\overline{x}$ is a global minimizer of $f$ on $S.$
On the other hand, since the mappings $f$ and $g$ are Lipschitz continuous around $\overline{x}$ and $\|w_{k}\|=1,$ it follows from Lemma 2.6 that the sequences $u_{k}$ and $v_{k}$ must be bounded. So we may assume that the sequences $u_{k},v_{k}$ and $w_{k}$ converge to vectors $\overline{u},\overline{v}$ and $\overline{w},$ respectively. Certainly, $\overline{u}\in\partial f(\overline{x}),$ $\overline{v}\in\partial\langle\overline{w},g\rangle(\overline{x})$ and $\overline{w}\in N_{C}(g(\overline{x}))$ with $\|\overline{w}\|=1$ (due to Lemmas 2.1, 2.3 and 4.1).
Observe that (5) can be rewritten as
$0$ $\in$ $\frac{1}{k}u_{k}+v_{k}+\frac{1}{k}\frac{\varepsilon_{k}}{\lambda_{k}}\mathbb{B}.$
Equivalently,
$\left\|\frac{1}{k}u_{k}+v_{k}\right\|$ $\leq$ $\frac{1}{k}\frac{\varepsilon_{k}}{\lambda_{k}}\ <\ \frac{2(f_{*}-\mathfrak{m})}{k(\|x_{k}\|+1)}.$
Letting $k\to\infty,$ one gets $\overline{v}=0,$ in contradiction to (A1).
Case 2: the sequence $\{x_{k}\}$ is unbounded.
By (4), the sequence $\{x_{k}^{\prime}\}$ is also unbounded. Passing to a subsequence if necessary, we can suppose that $\|x_{k}^{\prime}\|\to\infty$ as $k\to\infty.$ Note that (5) can be rewritten as
$0$ $\in$ $\frac{1}{{1+k}}u_{k}+\frac{k}{{1+k}}v_{k}+\frac{1}{{1+k}}\cdot\frac{\varepsilon_{k}}{\lambda_{k}}\mathbb{B}.$
By definition, then
$\nu(x_{k}^{\prime})$ $\leq$ $\left\|\frac{1}{{1+k}}u_{k}+\frac{k}{{1+k}}v_{k}\right\|\ \leq\ \frac{\varepsilon_{k}}{(1+k)\lambda_{k}}\ <\ \frac{2(f_{*}-\mathfrak{m})}{{(1+k)}(\|x_{k}\|+1)}.$
In combining this with (4), we obtain
$(\|x_{k}^{\prime}\|+1)\nu(x_{k}^{\prime})$ $<$ $\frac{(f_{*}-\mathfrak{m})}{3({1+k})}.$
Thus, $\|x_{k}^{\prime}\|\nu(x_{k}^{\prime})\to 0$ as $k\to\infty.$ On the other hand, we have shown that $x_{k}^{\prime}\not\in S,$ $\psi(x_{k}^{\prime})\to 0$ as $k\to\infty$ and the sequence $f(x_{k}^{\prime})$ has a cluster point, say, $t,$ which belongs to the interval $\mathfrak{m},f_{*}].$ Certainly $t\in K_{\infty}(f,S),$ in contradiction to (A2). The equivalence of (ii) with ($\star$), and hence with (i), is proved.
Finally, assume that the property (i) holds. By Theorem 3.1, one has for any $c>c_{*}$ that
$\mathrm{argmin}_{g(x)\in C}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big\{f(x)+c\,\mathrm{dist}(g(x),C)\big\}.$
Let $\overline{x}$ be a global minimum for the problem (P0). Then $\overline{x}$ is also a global minimum for the unconstrained optimization problem
$\textrm{minimize}_{x\in\mathbb{R}^{n}}\big\{f(x)+c\,\mathrm{dist}(g(x),C)\big\},$
and so $0\in\partial\big\{f(\cdot)+c\,\mathrm{dist}(g(\cdot),C)\big\}(\overline{x})$ due to Lemma 2.5. This, together with Lemmas 2.2, 2.4 and 2.7, yields
$0$ $\in$ $\partial f(\overline{x})+c\,\partial\langle w,g\rangle(\overline{x})$
for some $w\in N_{C}(g(\overline{x}))\cap\mathbb{B},$ as required. ∎
The assumptions (A1) and (A2) cannot be omitted as shown below.
Example 4.1.
(i) Let $n:=1,S:=\{x\in\mathbb{R}\mid x^{2}=0\}$ and $f(x):=x^{3}.$ Clearly, for any $c>0$ we have
$\inf_{x\in S}f(x)=0>-\infty=\inf_{x\in\mathbb{R}}\big\{f(x)+c\,\psi(x)\big\},$
where $\psi(x):=x^{2}.$ Observe that (A2) holds while (A1) does not.
(ii) Let $n:=2,S:=\{x:=(x_{1},x_{2})\in\mathbb{R}^{2}\ |\ x_{1}=0\}$ and $f(x):=e^{x_{1}x_{2}}.$ Clearly, for any $c>0$ we have
$\inf_{x\in S}f(x)=1>0=\inf_{x\in\mathbb{R}^{2}}\big\{f(x)+c\,\psi(x)\big\},$
where $\psi(x):=|x_{1}|.$ Observe that (A1) is satisfied, but (A2) is not.
(iii) Let $n:=2,S:=\{x:=(x_{1},x_{2})\in\mathbb{R}^{2}\mid x_{1}-x_{2}=0\}$ and $f(x):=x_{1}-x_{2}.$ Observe that the assumption (1.1) in [47] and the assumption (1.6) in [49] do not satisfy, and so we can not apply the results in these papers. On the other hand, it is easy to see that the assumptions (A1) and (A2) hold, and indeed, for any $c\geq 1$ we have
$\inf_{x\in S}f(x)=0=\inf_{x\in\mathbb{R}^{2}}\big\{f(x)+c\,\psi(x)\big\},$
where $\psi(x):=|x_{1}-x_{2}|.$
(iv) Let $n:=2,S:=\{x:=(x_{1},x_{2})\in\mathbb{R}^{2}\mid x_{1}=0\}$ and $f(x):=x_{1}^{3}+x_{2}^{2}.$ For any $c>0$ we have
$\inf_{x\in S}f(x)=0>-\infty=\inf_{x\in\mathbb{R}^{2}}\big\{f(x)+c\,\psi(x)\big\},$
where $\psi(x):=|x_{1}|.$ Observe that the assumptions (A1) and (A2) are satisfied.
Remark 4.3.
(i) Theorem 4.1 can be strengthened by replacing $f$ by $f_{0}+\delta_{X}$ with $f_{0}\colon\mathbb{R}^{n}\rightarrow\mathbb{R}$ being a locally Lipschitz function and $X$ being a closed subset of $\mathbb{R}^{n}.$ Since we do not use this fact, we leave the details for the interested reader.
(ii) The assumption (A2) holds trivially when $K_{\infty}(f,S)$ is an empty set. Consequently, Theorem 4.1 improves [47, Theorems 1.5 and 3.1] where it is assumed that $C=\{0\}\times\mathbb{R}_{-}^{m-p},$ $f$ and $g$ are locally Lipschitz, $f$ is coercive and the linear independence constraint qualification holds at every global minimum of $f$ on $S.$ In the finite-dimensional setting, Theorem 4.1 also improves [49, Theorems 1.3].
5. Exact penalty functions: the semi-algebraic case
In this section we investigate exact penalty functions for constrained optimization problems with semi-algebraic data. The main tool for this study will be Łojasiewicz inequalities on unbounded sets. Note that the results presented here are still valid for functions definable in a polynomially bounded o-minimal structure; however, to lighten the exposition, we do not pursue this extension here.
5.1. Łojasiewicz inequalities on unbounded sets
In this subsection, we present some Łojasiewicz inequalities on unbounded sets for (not necessarily continuous) semi-algebraic functions, and for the convenience of the reader, we provide detailed proofs to make the paper self-contained. The results presented here, which are inspired by the work of Dinh, Hà and Thao [16] (see also [13, 30, 32, 33]), will play a crucial role in establishing exact penalty properties in constrained optimization with semi-algebraic data.
Let us start with the following definition.
Definition 5.1.
Let $X$ be a subset of $\mathbb{R}^{n}$ and $\phi,\psi\colon X\rightarrow\mathbb{R}$ be functions. A sequence $\{x_{k}\}$ in $X$ is said to be
  1. (i)
    a sequence of the first type for the pair $(\phi,\psi)$ if $\psi(x_{k})\to 0$ and $|\phi(x_{k})|\geq\epsilon$ for some $\epsilon>0;$
  2. (ii)
    a sequence of the second type for the pair $(\phi,\psi)$ if the sequence $\psi(x_{k})$ is bounded and $|\phi(x_{k})|\to\infty.$
The following result improves [16, Proposition 3.7] in which $X=\mathbb{R}^{n},$ $\psi$ is the absolute value function of a polynomial and $\phi$ is the distance function $x\mapsto\mathrm{dist}(x,\psi^{-1}(0)).$
Lemma 5.1.
Let $X$ be a semi-algebraic subset of $\mathbb{R}^{n}$ and $\phi,\psi\colon X\rightarrow\mathbb{R}$ be nonnegative and semi-algebraic functions. The following two conditions are equivalent:
  1. (i)
    there exist some constants $c>0,\epsilon>0$ and $\alpha\in(0,1]$ such that
    $c\,[\psi(x)]^{\alpha}$ $\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in\psi^{-1}([0,\epsilon));$
  2. (ii)
    there is no sequence of the first type for the pair $(\phi,\psi).$
Proof.
(i) $\Rightarrow$ (ii) The implication is straightforward.
(ii) $\Rightarrow$ (i) Since there is no sequence of the first type for the pair $(\phi,\psi),$ we have
  1. Fact 1.
    $\phi(x)\to 0$ as $\psi(x)\to 0;$ in particular, $\phi(x)=0$ provided that $\psi(x)=0.$
On the other hand, in light of Theorem 2.1, $\psi(X)$ is a semi-algebraic set in $\mathbb{R},$ so it is a finite union of points and intervals (see Lemma 2.9). Consequently, there exists a constant $\epsilon>0$ such that either $(0,\epsilon)\cap\psi(X)=\emptyset$ or $(0,\epsilon)\subset\psi(X).$ In the first case, for all $x\in\psi^{-1}[0,\epsilon)$ we have $\psi(x)=0,$ and by Fact 1, $\phi(x)=0;$ consequently, (i) holds trivially for any choice of $c>0$ and $\alpha>0.$ So, assume that the second case holds, i.e., $(0,\epsilon)\subset\psi(X).$ Then for all $t\in(0,\epsilon),$ the set $\psi^{-1}(t)$ is nonempty and so the function
$\mu\colon(0,\epsilon)\to\mathbb{R}\cup\{\infty\},\quad t\mapsto\sup_{x\in\psi^{-1}(t)}\phi(x),$
is well defined. Observe that, by Theorem 2.1, the function $\mu$ is semi-algebraic. Moreover, it follows from Fact 1 that $\mu(t)\to 0$ as $t\to 0^{+}.$ By Lemma 2.12 and decreasing $\epsilon$ if necessary, we can see that the function $\mu$ is finite and either constant or strictly monotone. There are two cases to consider.
Case 1.1:
$\mu$ is constant.
Then $\mu(t)=0$ for all $t\in(0,\epsilon).$ It follows that for all $x\in\psi^{-1}(0,\epsilon),$
$\psi(x)\geq 0=\phi(x).$
Then (i) holds trivially for any choice of $c>0$ and $\alpha>0.$
Case 1.2:
$\mu$ is not constant.
Thanks to Lemma 2.11(i), we can write
$\mu(t)=a\,t^{\alpha}+o(t^{\alpha})\quad\textrm{ as }\quad t\to 0^{+}$
for some constants $a>0$ and $\alpha\in\mathbb{Q}.$ Observe that $\alpha>0$ because $\mu(t)\to 0$ as $t\to 0^{+}.$ Reducing $\epsilon$ and replacing $\alpha$ by $\min\{\alpha,1\}$ (if necessary), we can see that
$\mu(t)\leq 2a\,t^{\alpha}\quad\textrm{ for }\quad t\in(0,\epsilon).$
Hence (i) holds when $c:=2a>0.$
The following corollary improves [32, Lemma 3.2] in which $X$ and $\phi$ are assumed to be compact and continuous, respectively; see also [30, Theorem 2.10].
Corollary 5.1.
Let $X$ be a semi-algebraic subset of $\mathbb{R}^{n}$ and $\phi,\psi\colon X\rightarrow\mathbb{R}$ be nonnegative and semi-algebraic functions. If in addition $\phi$ is bounded from above, then the following two conditions are equivalent:
  1. (i)
    there exist some constants $c>0$ and $\alpha\in(0,1]$ such that
    $c\,[\psi(x)]^{\alpha}$ $\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in X;$
  2. (ii)
    there is no sequence of the first type for the pair $(\phi,\psi).$
Proof.
(i) $\Rightarrow$ (ii) The implication is straightforward.
(ii) $\Rightarrow$ (i) By Lemma 5.1, we can find constants $c_{1}>0,\epsilon>0$ and $\alpha\in(0,1]$ such that
$c_{1}[\psi(x)]^{\alpha}$ $\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in\psi^{-1}([0,\epsilon)).$
On the other hand, since $\phi$ is bounded from above, there exists a constant $M>0$ such that $\phi(x)\leq M$ for all $x\in X.$ Hence for all $x\in\psi^{-1}([\epsilon,\infty))$ we have
$[\psi(x)]^{\alpha}$ $\geq$ $\epsilon^{\alpha}\ =\ \frac{\epsilon^{\alpha}}{M}M\ \geq\ \frac{\epsilon^{\alpha}}{M}\,\phi(x).$
Letting $c:=\max\{c_{1},\frac{M}{\epsilon^{\alpha}}\},$ we get the desired conclusion. ∎
Note here that the assumption of boundedness of the function $\phi$ cannot be omitted; see Example 5.1.
Corollary 5.2.
Let $A,B$ be closed semi-algebraic subsets of $\mathbb{R}^{n}$ such that $A\cap B\neq\emptyset.$ Then there exist some constants $c>0$ and $\alpha\in(0,1]$ such that
$c\,\big(\mathrm{dist}(x,A)+\mathrm{dist}(x,B)\big)^{\alpha}$ $\geq$ $\frac{\mathrm{dist}(x,A\cap B)}{1+\|x\|^{2}}\quad\textrm{ for all }\quad x\in\mathbb{R}^{n}.$
Proof.
Define the functions $\phi,\psi\colon\mathbb{R}^{n}\to\mathbb{R}$ by
$\phi(x):=\frac{\mathrm{dist}(x,A\cap B)}{1+\|x\|^{2}}\quad$ and $\quad\psi(x):=\mathrm{dist}(x,A)+\mathrm{dist}(x,B),$
which are nonnegative, continuous and semi-algebraic (by Theorem 2.1). Note that $\phi(x)=0$ when $\psi(x)=0$ and that $\phi(x)\to 0$ as $\|x\|\to\infty.$ Thus, the function $\phi$ is bounded from above and there is no sequence of the first type for the pair $(\phi,\psi).$ The desired conclusion follows directly from Corollary 5.1. ∎
Lemma 5.2.
Let $X$ be a semi-algebraic subset of $\mathbb{R}^{n}$ and $\phi,\psi\colon X\rightarrow\mathbb{R}$ be nonnegative and semi-algebraic functions. The following two conditions are equivalent:
  1. (i)
    there exist some constants $c>0,R>0$ and $\beta\geq 1$ such that
    $c\,[\psi(x)]^{\beta}$ $\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in\psi^{-1}((R,\infty));$
  2. (ii)
    there is no sequence of the second type for the pair $(\phi,\psi).$
Proof.
(i) $\Rightarrow$ (ii) The implication is straightforward.
(ii) $\Rightarrow$ (i) Since there is no sequence of the second type, we have
  1. Fact 2.
    For any $a,b\in\mathbb{R}$ with $0\leq a\leq b<\infty,$ there exists a constant $M>0$ such that $\phi(x)\leq M$ for all $x\in\psi^{-1}([a,b]).$
On the other hand, in light of Theorem 2.1, $\psi(X)$ is a semi-algebraic set in $\mathbb{R},$ so it is a finite union of points and intervals (see Lemma 2.9). Consequently, there exists a constant $R>0$ such that either $(R,\infty)\cap\psi(X)=\emptyset$ or $(R,\infty)\subset\psi(X).$ If the first case happens, then (i) holds trivially for any choice of $c>0$ and $\beta>0$ and there is nothing to prove. Hence, assume that $(R,\infty)\subset\psi(X).$ Then for all $t\in(R,\infty),$ the set $\psi^{-1}(t)$ is nonempty. It follows from Fact 2 that the function
$\mu\colon(R,\infty)\to\mathbb{R},\quad t\mapsto\sup_{x\in\psi^{-1}(t)}\phi(x),$
is well defined. Observe that, by Theorem 2.1, the function $\mu$ is semi-algebraic. By Lemma 2.12 and increasing $R$ if necessary, we can see that the function $\mu$ is either constant or strictly monotone. There are two cases to consider.
Case 2.1:
$\mu$ is constant, say, $c^{\prime}.$
Then $c^{\prime}\geq 0$ and for all $x\in\psi^{-1}(R,\infty)$ we have
$\psi(x)\geq R=\frac{R}{c^{\prime}+1}(c^{\prime}+1)\geq\frac{R}{c^{\prime}+1}\phi(x).$
Then (i) holds trivially when $c:=\frac{c^{\prime}+1}{R}$ and $\beta:=1.$
Case 2.2:
$\mu$ is not constant.
Thanks to Lemma 2.11(ii), we can write
$\mu(t)=b\,t^{\beta}+o(t^{\beta})\quad\textrm{ as }\quad t\to\infty$
for some constants $b>0$ and $\beta\in\mathbb{Q}.$ Increasing $R$ and replacing $\beta$ by $\max\{\beta,1\}$ (if necessary), we can see that
$\mu(t)\leq 2b\,t^{\beta}\quad\textrm{ for }\quad t\in(R,\infty).$
Hence (i) holds when $c:=2b>0.$
The following result improves [16, Proposition 3.10] in which $X=\mathbb{R}^{n},$ $\psi$ is the absolute value function of a polynomial and $\phi$ is the distance function $x\mapsto\mathrm{dist}(x,\psi^{-1}(0)).$
Proposition 5.1.
Let $X$ be a semi-algebraic subset of $\mathbb{R}^{n}$ and $\phi,\psi\colon X\rightarrow\mathbb{R}$ be nonnegative and semi-algebraic functions. The following two conditions are equivalent:
  1. (i)
    there exist some constants $c>0,\alpha>0,$ and $\beta>0$ with $\alpha\leq 1\leq\beta$ such that
    $c\left([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\right)$ $\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in X;$
  2. (ii)
    there are no sequences of the first and second types for the pair $(\phi,\psi).$
Proof.
(i) $\Rightarrow$ (ii) The implication is straightforward.
(ii) $\Rightarrow$ (i) By Lemma 5.1, there exist some constants $c_{1}>0,\epsilon>0$ and $\alpha\in(0,1]$ such that
$c_{1}\,[\psi(x)]^{\alpha}$ $\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in\psi^{-1}([0,\epsilon)).$
In view of Lemma 5.2, there exist some constants $c_{2}>0,R>\epsilon$ and $\beta\geq 1$ such that
$c_{2}\,[\psi(x)]^{\beta}$ $\geq$ $\phi(x)\quad\textrm{ for all }\quad x\in\psi^{-1}((R,\infty)).$
On the other hand, it follows from Fact 2 in the proof of Lemma 5.2 that there exists a constant $M>0$ such that for all $x\in\psi^{-1}([\epsilon,R])$ we have $\phi(x)\leq M,$ and hence
$[\psi(x)]^{\alpha}+[\psi(x)]^{\beta}$ $\geq$ $\epsilon^{\alpha}+\epsilon^{\beta}\ =\ \frac{\epsilon^{\alpha}+\epsilon^{\beta}}{M}M\ \geq\ \frac{\epsilon^{\alpha}+\epsilon^{\beta}}{M}\,\phi(x).$
Letting $c:=\max\{c_{1},c_{2},\frac{M}{\epsilon^{\alpha}+\epsilon^{\beta}}\},$ we get the desired conclusion. ∎
The following simple example shows that the exponents $\alpha$ and $\beta$ in Proposition 5.1 are different in general.
Example 5.1.
Consider the polynomial functions $\phi(x_{1},x_{2}):=x_{1}^{2}+x_{2}^{2}$ and $\psi(x_{1},x_{2}):=x_{1}^{2}+x_{2}^{4}$ on $\mathbb{R}^{2}.$ It is not hard to see that there are no constants $c>0$ and $\alpha>0$ such that
$c\,[\psi(x_{1},x_{2})]^{\alpha}$ $\geq$ $\phi(x_{1},x_{2})\quad\textrm{ for all }\quad(x_{1},x_{2})\in\mathbb{R}^{2}.$
On the other hand, it holds that
$[\psi(x_{1},x_{2})]^{\frac{1}{2}}+\psi(x_{1},x_{2})$ $\geq$ $\phi(x_{1},x_{2})\quad\textrm{ for all }\quad(x_{1},x_{2})\in\mathbb{R}^{2}.$
We close this subsection by presenting a necessary and sufficient condition for a semi-algebraic function to have a global Hölderian error bound; this result improves [25, Theorem 3.4], where the continuity of the function in question is assumed.
Corollary 5.3.
Let $f\colon\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a semi-algebraic function, and $S:=\{x\in\mathbb{R}^{n}\mid f(x)\leq 0\}\neq\emptyset.$ The following two conditions are equivalent:
  1. (i)
    there exist some constants $c>0,\alpha>0,$ and $\beta>0$ with $\alpha\leq 1\leq\beta$ such that
    $c\left([f(x)]_{+}^{\alpha}+[f(x)]_{+}^{\beta}\right)$ $\geq$ $\mathrm{dist}(x,S),\quad\textrm{ for all }\quad x\in\mathbb{R}^{n};$
  2. (ii)
    for any sequence of points $x_{k}$ in $\mathbb{R}^{n}$ one has
    1. (ii1)
      if $f(x_{k})\to 0,$ then $\mathrm{dist}(x,S)\to 0;$
    2. (ii2)
      if $\mathrm{dist}(x,S)\to\infty,$ then $f(x_{k})\to\infty.$
Proof.
This applies Proposition 5.1 to the semi-algebraic functions $\psi(x):=[f(x)]_{+}$ and $\phi(x):=\mathrm{dist}(x,S)$ for $x\in\mathbb{R}^{n}.$
5.2. Necessary and sufficient conditions for exact penalization
The results presented in this subsection are inspired by the works of Warga [44, Theorem 1], Dedieu [12, Theorem 3.1] and Luo, Pang and Ralph [34, Chapter 2]. Indeed, consider the problem (P) with $S$ being a nonempty semi-algebraic subset of $\mathbb{R}^{n}$ and $f\colon\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ being a semi-algebraic function. Assume that the optimal value $f_{*}:=\inf_{S}f$ is finite and let $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ be a residual function for $S$ which is semi-algebraic. Let’s start with the following.
Remark 5.1.
If for some $c_{*}>0$ we have
$\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c_{*}\,\psi(x)\big\},$
then by Theorem 3.1, for any sequence of points $x_{k}$ in $\mathbb{R}^{n}$ with $\psi(x_{k})\to 0$ the sequence $f_{*}-f(x_{k})]_{+}$ must converge to $0.$ The converse does not hold in general; see Example 5.2(ii) below. On the other hand, we have the following result.
Theorem 5.1.
Assume that for any sequence of points $x_{k}$ in $\mathbb{R}^{n}$ with $\psi(x_{k})\to 0$ one has $f_{*}-f(x_{k})]_{+}\to 0.$ Then there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that for all $c>c_{*},$
$\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c\,\big(1+[f(x)]^{2}\big)[\psi(x)]^{\alpha}\big\},$
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big\{f(x)+c\,\big(1+[f(x)]^{2}\big)[\psi(x)]^{\alpha}\big\}.$
Proof.
Let $A:=(-\infty,f_{*}]\times\{0\}$ and $B:=\mathrm{cl}(f,\psi)(\mathrm{dom}f).$ Then $A$ and $B$ are closed semi-algebraic sets in $\mathbb{R}^{2}.$ Moreover, our assumption yields $A\cap B=\{(f_{*},0)\}.$ In view of Corollary 5.2, there are some constants $c_{*}>0$ and $\alpha\in(0,1]$ such that
$c_{*}\big(\mathrm{dist}(y,A)+\mathrm{dist}(y,B)\big)^{\alpha}$ $\geq$ $\frac{\mathrm{dist}(y,A\cap B)}{1+\|y\|^{2}}\quad\textrm{ for all }\quad y\in\mathbb{R}^{2}.$
Take $x\in\mathbb{R}^{n}$ with $f(x)<f_{*}$ and let $y:=(f(x),0).$ We have
$\mathrm{dist}(y,A)$ $=$ $0,$
$\mathrm{dist}(y,B)$ $\leq$ $\|(f(x),0)-(f(x),\psi(x))\|=\psi(x),$
$\mathrm{dist}(y,A\cap B)$ $=$ $\|(f(x),0)-(f_{*},0)\|=f_{*}-f(x).$
Hence
$c_{*}[\psi(x)]^{\alpha}$ $\geq$ $\frac{f_{*}-f(x)}{1+[f(x)]^{2}},$
or equivalently,
$c_{*}{\big(1+[f(x)]^{2}\big)}[\psi(x)]^{\alpha}$ $\geq$ ${f_{*}-f(x)}.$
Clearly, this inequality is true for any $x\in\mathbb{R}^{n}$ with $f(x)\geq f_{*}.$ By Theorem 3.1, we get the desired conclusion. ∎
Remark 5.2.
(i) Theorem 5.1 can fail when “semi-algebraic" is replaced by “subanalytic” (cf. [4] for this notion); see Example 5.3.
(ii) In general, the converse of Theorem 5.1 does not hold and the term $(1+[f(x)]^{2})$ cannot be omitted; see the next example.
Example 5.2.
(i) Let $n:=1,S:=\{0\}$ and consider the semi-algebraic functions
$f(x)$ $:=$ $x\quad\textrm{ and }\quad\psi(x)\ :=\ \frac{|x|}{1+x^{2}}.$
Then $\psi$ is a residual function for $S$ and it holds that
$\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}}\big\{f(x)+(1+[f(x)]^{2})[\psi(x)]\big\},$
while
$\psi(x)\to 0\quad\textrm{ and }\quad[f_{*}-f(x)]_{+}\to\infty\quad\textrm{ as }\ x\to-\infty.$
(ii) Let $n:=1,S:=\{0\}$ and $f(x):=x.$ Consider the semi-algebraic function $\psi\colon\mathbb{R}\to\mathbb{R}_{+}$ defined by
$\psi(x)$ $:=$ $\begin{cases}|x|&\textrm{ if }|x|\leq 1,\\ 1&\textrm{ otherwise,}\end{cases}$
which is a residual function for $S.$ A direct calculation shows that
$\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}}\big\{f(x)+(1+[f(x)]^{2})[\psi(x)]\big\},$
while for any $c>0$ and any $\alpha>0,$
$\inf_{x\in S}f(x)$ $=$ $0\ >\ -\infty\ =\ \inf_{x\in\mathbb{R}}\big\{f(x)+c\,[\psi(x)]^{\alpha}\big\}.$
The following result shows that the term $(1+[f(x)]^{2})$ can be removed provided that the function $f$ is bounded from below on $\mathbb{R}^{n}.$
Theorem 5.2.
If $f$ is bounded from below on $\mathbb{R}^{n},$ then the following properties are equivalent:
  1. (i)
    there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that for all $c>c_{*},$
    $\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c\,[\psi(x)]^{\alpha}\big\};$
  2. (ii)
    there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that
    $c_{*}[\psi(x)]^{\alpha}$ $\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad x\in\mathbb{R}^{n};$
  3. (iii)
    for any sequence of points $x_{k}$ in $\mathbb{R}^{n}$ with $\psi(x_{k})\to 0,$ one has $f_{*}-f(x_{k})]_{+}\to 0.$
When these equivalent properties hold, one has moreover for all $c>c_{*}$ that
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big\{f(x)+c\,[\psi(x)]^{\alpha}\big\}.$
Proof.
The equivalence of (i) and (ii) as well as the last conclusion follow from Theorem 3.1. To get the equivalence of (ii) and (iii), we can apply Corollary 5.1 to the functions $\phi$ and $\psi$ with $\phi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ being defined by
$\phi(x)$ $:=$ $[f_{*}-f(x)]_{+}\quad\textrm{ for }\quad x\in\mathbb{R}^{n},$
which is bounded from above on $\mathbb{R}^{n}.$
Remark 5.3.
(i) In general, the assumption in Theorem 5.2 that $f$ is bounded from below on $\mathbb{R}^{n}$ cannot be dropped; see Example 5.2(ii). On the other hand, we have Theorem 5.3 below.
(ii) A semi-algebraic version of [44, Theorem 1] as well as [12, Theorem 3.1] is followed directly from Theorem 5.2. To see this, let $f:=f_{0}+\delta_{X},$ $f_{0}\colon X\to\mathbb{R}$ be continuous and semi-algebraic, $X\subset\mathbb{R}^{n}$ and $S\subset X$ nonempty, compact and semi-algebraic. Also, let $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ be a residual function for $S$ which is continuous and semi-algebraic. Then it is not hard to see that $f$ is bounded from below on $\mathbb{R}^{n}$ and that the condition (iii) of Theorem 5.2 is satisfied, and so there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that for all $c>c_{*},$
$\inf_{x\in S}f_{0}(x)$ $=$ $\inf_{x\in X}\big\{f_{0}(x)+c\,[\psi(x)]^{\alpha}\big\},$
$\mathrm{argmin}_{S}f_{0}(x)$ $=$ $\mathrm{argmin}_{X}\big\{f_{0}(x)+c\,[\psi(x)]^{\alpha}\big\}.$
It is worth noting here that, in the subanalytic setting, these relations are still valid (see [12, Theorem 3.1]) but can fail when $X$ is unbounded. Also, the equivalence of (iii) with (ii) (and hence with (i)) can fail when “semi-algebraic" is replaced by “subanalytic”; see the next example.
Example 5.3.
Let $n:=1,$ $S:=\{0\}$ and consider the continuous subanalytic functions
$f(x)$ $:=$ $\begin{cases}1-\frac{1}{x}&\textrm{ if }x\geq 1,\\ -x+1&\textrm{ if }0\leq x<1,\\ 1&\textrm{ otherwise,}\end{cases}$
and
$\psi(x)$ $:=$ $\begin{cases}e^{-x+1}&\textrm{ if }x\geq 1,\\ |x|&\textrm{ otherwise.}\end{cases}$
Certainly $f_{*}:=\inf_{x\in S}f(x)=1$ and $\psi$ is a residual function for $S.$ Moreover, for any sequence of points $x_{k}$ in $\mathbb{R}$ with $\psi(x_{k})\to 0,$ one has $f_{*}-f(x_{k})]_{+}\to 0.$ However, the equivalent properties (i) and (ii) in Theorem 5.2 fail to hold. This example also shows that [44, Theorem 1] and [12, Theorem 3.1] can fail for optimization problems over unbounded sets.
Theorem 5.3.
The following properties are equivalent:
  1. (i)
    there exist constants $c_{*}>0,\alpha>0,$ and $\beta>0$ with $\alpha\leq 1\leq\beta$ such that for all $c>c_{*},$
    $\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\left\{f(x)+c\left([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\right)\right\};$
  2. (ii)
    there exist constants $c_{*}>0,\alpha>0,$ and $\beta>0$ with $\alpha\leq 1\leq\beta$ such that
    $c_{*}\left([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\right)$ $\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad x\in\mathbb{R}^{n};$
  3. (iii)
    for any sequence of points $x_{k}$ in $\mathbb{R}^{n}$ one has
    1. (iii1)
      if $\psi(x_{k})\to 0,$ then $f_{*}-f(x_{k})]_{+}\to 0;$
    2. (iii2)
      if $f(x_{k})\to-\infty,$ then $\psi(x_{k})\to\infty.$
When these equivalent properties hold, one has moreover for all $c>c_{*}$ that
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\left\{f(x)+c\left([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\right)\right\}.$
Proof.
The equivalence of (i) and (ii) as well as the last conclusion follow from Theorem 3.1, while the equivalence of (ii) and (iii) is known from Proposition 5.1. ∎
Remark 5.4.
(i) The exponents $\alpha$ and $\beta$ can take, respectively, to be $\frac{1}{N}$ and $N$ for some integer $N>0.$
(ii) To incorporate an abstract constraint of the form $x\in X$ with $X$ being a semi-algebraic subset of $\mathbb{R}^{n},$ we simply replace $f$ by $f+\delta_{X}.$
(iii) The term $\psi(x)]^{\beta}$ cannot be omitted as shown in the following example.
Example 5.4.
Let $n:=2$ and consider the polynomial functions $f(x_{1},x_{2}):=-x_{1}^{2}-x_{2}^{2}$ and $\psi(x_{1},x_{2}):=x_{1}^{2}+x_{2}^{4}.$ By definition, we have that $\psi$ is a residual function for $S:=\{(0,0)\}$ and $f_{*}:=\inf_{(x_{1},x_{2})\in S}f(x_{1},x_{2})=0.$ Moreover, it is not hard to see that there are no constants $c>0$ and $\alpha>0$ such that
$c[\psi(x_{1},x_{2})]^{\alpha}$ $\geq$ $[f_{*}-f(x_{1},x_{2})]_{+}\quad\textrm{ for all }\quad(x_{1},x_{2})\in\mathbb{R}^{2}.$
On the other hand, it holds that
$[\psi(x_{1},x_{2})]^{\frac{1}{2}}+\psi(x_{1},x_{2})$ $\geq$ $[f_{*}-f(x_{1},x_{2})]_{+}\quad\textrm{ for all }\quad(x_{1},x_{2})\in\mathbb{R}^{2},$
which, together with Theorem 5.3, implies that for all $c>1,$
$\inf_{(x_{1},x_{2})\in S}f(x_{1},x_{2})$ $=$ $\inf_{(x_{1},x_{2})\in\mathbb{R}^{2}}\left\{f(x_{1},x_{2})+c\left([\psi(x_{1},x_{2})]^{\frac{1}{2}}+\psi(x_{1},x_{2})\right)\right\}.$
In the rest of this subsection, we consider the problem (P) with $f:=f_{0}+\delta_{X},$ $f_{0}\colon X\to\mathbb{R}$ (not necessarily semi-algebraic) Lipschitz, $X\subset\mathbb{R}^{n}$ and $S\subset X$ nonempty and semi-algebraic. Let $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ be a residual function for $S$ which is semi-algebraic.
Theorem 5.4.
Assume that for any sequence of points $x_{k}$ in $X,$
  1. (A1)
    if $\psi(x_{k})\to 0$ with $x_{k}\in X,$ then $\mathrm{dist}(x_{k},S)\to 0;$
  2. (A2)
    the distance function $\mathrm{dist}(\cdot,S)$ is bounded from above on $X.$
Then there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that for all $c>c_{*},$
$\inf_{x\in S}f_{0}(x)$ $=$ $\inf_{x\in X}\big\{f_{0}(x)+c\,[\psi(x)]^{\alpha}\big\},$
$\mathrm{argmin}_{S}f_{0}(x)$ $=$ $\mathrm{argmin}_{X}\big\{f_{0}(x)+c\,[\psi(x)]^{\alpha}\big\}.$
Proof.
By Theorem 3.1, it suffices to show the first equality.
According to Corollary 5.1, there exist some constants $c_{1}>0$ and $\alpha\in(0,1]$ such that
$c_{1}\,[\psi(x)]^{\alpha}$ $\geq$ $\mathrm{dist}(x,S)\quad\textrm{ for all }\quad x\in X.$
Let $L>0$ be the Lipschitzian modulus of $f_{0}$ on $X,$ i.e., for all $x,x^{\prime}\in X,$
$|f_{0}(x)-f_{0}(x^{\prime})|$ $\leq$ $L\|x-x^{\prime}\|.$
Also let $c_{*}$ be an arbitrary scalar greater than $c_{1}L.$ Let $c>c_{*}$ and $x\in X$ an arbitrary point. Take any $\epsilon>0.$ By definition, there exists $x^{\prime}\in S$ such that
$\|x-x^{\prime}\|$ $\leq$ $\mathrm{dist}(x,S)+\epsilon.$
On the other hand, we have $f_{0}(x^{\prime})\geq\inf_{S}f_{0}$ and $f_{0}(x)-f_{0}(x^{\prime})\geq-L\|x-x^{\prime}\|.$ Thus,
$f_{0}(x)+c\,[\psi(x)]^{\alpha}$ $=$ $f_{0}(x^{\prime})+\big(f_{0}(x)-f_{0}(x^{\prime})\big)+c\,[\psi(x)]^{\alpha}$
$\geq$ $f_{0}(x^{\prime})-L\|x-x^{\prime}\|+\frac{c}{c_{1}}\,\mathrm{dist}(x,S)$
$\geq$ $f_{0}(x^{\prime})-L\|x-x^{\prime}\|+\frac{c}{c_{1}}\big(\|x-x^{\prime}\|-\epsilon\big)$
$=$ $f_{0}(x^{\prime})+\left(\frac{c}{c_{1}}-L\right)\|x-x^{\prime}\|-\frac{c}{c_{1}}\epsilon$
$\geq$ $\inf_{S}f_{0}-\frac{c}{c_{1}}\epsilon.$
Taking the limit as $\epsilon\searrow 0$ yields the inequality
$f_{0}(x)+c\,[\psi(x)]^{\alpha}$ $\geq$ $\inf_{S}f_{0}.$
Since $x$ was arbitrary in $X,$ we get
$\inf_{x\in X}\left\{f_{0}(x)+c\,[\psi(x)]^{\alpha}\right\}$ $\geq$ $\inf_{x\in S}f_{0}(x).$
The opposite inequality holds trivially, so the desired equality follows. ∎
Remark 5.5.
(i) A semi-algebraic version of [34, Theorem 2.1.2] is derived directly from Theorem 5.4 because the conditions of the theorem hold automatically when $X$ is compact and $\psi$ is continuous.
(ii) The conclusions of Theorem 5.4 can fail when “semi-algebraic" is replaced by “subanalytic”; see the next example.
Example 5.5.
Let $n:=2,$ $X:=\left\{x:=(x_{1},x_{2})\in\mathbb{R}^{2}\mid|x_{2}|\leq 1\right\},$ $S:=\left\{(x_{1},x_{2})\in\mathbb{R}^{2}\mid x_{2}=0\right\}$ and consider the subanalytic functions $f_{0}(x):=-|x_{2}|$ and
$\psi(x)$ $:=$ $\begin{cases}e^{-\frac{1}{x_{2}^{2}}}&\textrm{ if }x_{2}\neq 0,\\ 0&\textrm{ otherwise.}\end{cases}$
Clearly, ${\rm dist}(x,S)=|x_{2}|\leq 1$ for all $x\in X$ and $f_{*}:=\inf_{x\in S}f_{0}(x)=0.$ Moreover, it is easy to check that for any sequence of points $x_{k}\in X$ with $\psi(x_{k})\to 0,$ one has $\mathrm{dist}(x_{k},S)\to 0.$ However, the conclusions of Theorem 5.4 are not valid. This example also shows that [34, Theorem 2.1.2] can fail for optimization problems over unbounded sets.
Theorem 5.5.
Assume that for any sequence of points $x_{k}$ in $X$ the following properties hold
  1. (A1)
    if $\psi(x_{k})\to 0,$ then $\mathrm{dist}(x_{k},S)\to 0;$
  2. (A2)
    if $\mathrm{dist}(x_{k},S)\to\infty,$ then $\psi(x_{k})\to\infty.$
Then there exist constants $c_{*}>0,\alpha>0,$ and $\beta>0$ with $\alpha\leq 1\leq\beta$ such that for all $c>c_{*},$
$\inf_{x\in S}f_{0}(x)$ $=$ $\inf_{x\in X}\left\{f_{0}(x)+c\big([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\big)\right\},$
$\mathrm{argmin}_{S}f_{0}(x)$ $=$ $\mathrm{argmin}_{X}\left\{f_{0}(x)+c\big([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\big)\right\}.$
Proof.
The proof is similar to the one of Theorem 5.4 with taking into account Proposition 5.1 instead of Corollary 5.1. ∎
Example 5.6.
Let $f_{0}\colon\mathbb{R}^{n}\to\mathbb{R}$ be a Lipschitz function and
$S$ $:=$ $\{x\in\mathbb{R}^{n}\mid g_{1}(x)\leq 0,\ldots,g_{m}(x)\leq 0\},$
where $(g_{1},\ldots,g_{m})\colon\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a convenient and non-degenerate at infinity polynomial mapping (see Section 6 below). Certainly the semi-algebraic function
$\psi\colon\mathbb{R}^{n}\to\mathbb{R},\quad x\mapsto\max\{g_{1}(x),\ldots,g_{m}(x),0\},$
is residual for $S.$ Moreover, in view of [25, Theorem 3.7], there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that
$c_{*}\,\big([\psi(x)]^{\alpha}+[\psi(x)]\big)$ $\geq$ $\mathrm{dist}(x,S)\quad\textrm{ for all }\quad x\in\mathbb{R}^{n};$
in particular, the properties (A1) and (A2) in Theorem 5.5 are satisfied. By a similar argument to the one given in the proof of Theorem 5.4, we have for all $c>c_{*}$ that
$\inf_{x\in S}f_{0}(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\left\{f_{0}(x)+c\big([\psi(x)]^{\alpha}+[\psi(x)]\big)\right\},$
$\mathrm{argmin}_{S}f_{0}(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\left\{f_{0}(x)+c\big([\psi(x)]^{\alpha}+[\psi(x)]\big)\right\}$
(perhaps after increasing $c_{*}$).
6. Exact penalty functions: the non-degenerate polynomial case
In this section, we provide a necessary and sufficient condition for the exactness of a penalty function for non-degenerate polynomial optimization problems. This result involves the theory of Newton polyhedra. Let us start with some notation and definitions.
6.1. Newton polyhedra and non-degeneracy conditions
Given a nonempty set $J\subset\{1,\ldots,n\},$ we define
$\mathbb{R}^{J}$ $:=$ $\{x\in\mathbb{R}^{n}\mid x_{j}=0,\textrm{ for all }j\not\in J\}.$
We denote by $\mathbb{Z}_{+}$ the set of non-negative integers. If $\kappa=(\kappa_{1},\ldots,\kappa_{n})\in\mathbb{Z}_{+}^{n},$ we denote by $x^{\kappa}$ the monomial $x_{1}^{\kappa_{1}}\cdots x_{n}^{\kappa_{n}}.$
A subset $\Gamma\subset{\mathbb{R}}^{n}_{+}$ is a Newton polyhedron if there exists a finite subset $A\subset{\mathbb{Z}}^{n}_{+}$ such that $\Gamma$ is the convex hull in ${\mathbb{R}}^{n}$ of $A.$ We say that $\Gamma$ is the Newton polyhedron determined by $A$ and write $\Gamma=\Gamma(A).$ A Newton polyhedron $\Gamma$ is convenient if it intersects each coordinate axis at a point different from the origin 0 in $\mathbb{R}^{n}.$
Given a Newton polyhedron $\Gamma$ and a vector $q\in{\mathbb{R}}^{n},$ we define
$d(q,\Gamma)$ $:=$ $\min\{\langle q,\kappa\rangle\mid\kappa\in\Gamma\},$
$\Delta(q,\Gamma)$ $:=$ $\{\kappa\in\Gamma\mid\langle q,\kappa\rangle=d(q,\Gamma)\}.$
By definition, for each nonzero vector $q\in\mathbb{R}^{n},$ $\Delta(q,\Gamma)$ is a closed face of $\Gamma.$ Conversely, if $\Delta$ is a closed face of $\Gamma$, then there exists a nonzero vector $q\in\mathbb{R}^{n}$ such that $\Delta=\Delta(q,\Gamma),$ where we can in fact assume that $q\in\mathbb{Q}^{n}$ since $\Gamma$ is an integer polyhedron. The dimension of a face $\Delta$ is the minimum of the dimensions of the affine subspaces containing $\Delta.$ The faces of $\Gamma$ of dimension 0 are the vertices of $\Gamma.$
Let $f\colon\mathbb{R}^{n}\to\mathbb{R}$ be a polynomial function. Suppose that $f$ is written as $f=\sum_{\kappa}c_{\kappa}x^{\kappa}.$ The support of $f,$ denoted by $\mathrm{supp}(f),$ is the set of $\kappa\in\mathbb{Z}_{+}^{n}$ such that $c_{\kappa}\neq 0.$ The Newton polyhedron (at infinity) of $f$, denoted by $\Gamma(f),$ is the convex hull in $\mathbb{R}^{n}$ of the set $\mathrm{supp}(f),$ i.e., $\Gamma(f)=\Gamma(\mathrm{supp}(f)).$ The polynomial $f$ is convenient if $\Gamma(f)$ is convenient. For each (closed) face $\Delta$ of $\Gamma(f),$ we will denote
$f_{\Delta}(x)$ $:=$ $\sum_{\kappa\in\Delta}c_{\kappa}x^{\kappa}.$
Remark 6.1.
Let $\Delta:=\Delta(q,\Gamma(f))$ for some nonzero vector $q:=(q_{1},\ldots,q_{n})\in\mathbb{R}^{n}.$ By definition, $f_{\Delta}$ is a weighted homogeneous polynomial of type $(q,d:=d(q,\Gamma(f))),$ i.e., we have for all $t>0$ and all $x\in\mathbb{R}^{n},$
$f_{\Delta}(t^{q_{1}}x_{1},\ldots,t^{q_{n}}x_{n})$ $=$ $t^{d}f_{\Delta}(x_{1},\ldots,x_{n}).$
This implies the Euler relation
$\sum_{j=1}^{n}q_{j}x_{j}\frac{\partial f_{\Delta}}{\partial x_{j}}(x)$ $=$ $d\cdot f_{\Delta}(x).$ (6)
Definition 6.1.
Let $F:=(f_{1},\ldots,f_{m})\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m},1\leq m\leq n,$ be a polynomial mapping.
  1. (i)
    The mapping $F$ is convenient if its component polynomials are convenient.
  2. (ii)
    The mapping $F$ is Khovanskii non-degenerate at infinity if for any vector [^1] The number of vectors $q\in\mathbb{R}^{n}$ is infinite; however, there exists a finite number of faces $\Delta_{i}$ of the Newton polyhedron $\Gamma(f_{i}).$ $q\in\mathbb{R}^{n}$ with $d(q,\Gamma(f_{i}))<0$ for $i=1,\ldots,m,$ the semi-algebraic set
    $\{x\in(\mathbb{R}^{*})^{n}\mid f_{i,\Delta_{i}}(x)=0\textrm{ for }i=1,\ldots,m\},$
    is a complete intersection singularity, that is there are no $x\in(\mathbb{R}^{*})^{n}$ and $\lambda\in(\mathbb{R}^{*})^{m}$ such that
[…]
    where $\Delta_{i}:=\Delta(q,\Gamma(f_{i}))$ and $\mathbb{R}^{*}:=\mathbb{R}\setminus\{0\}.$
  3. (iii)
    The mapping $F$ is non-degenerate at infinity if for each $k$-tuple $(i_{1},\ldots,i_{k})$ of integers with $1\leq i_{1}<\cdots<i_{k}\leq m,$ the polynomial mapping R^n →R^k, x ↦(f_i_1(x), …, f_i_k(x)), is Khovanskii non-degenerate at infinity.
Remark 6.2.
(i) The main result in this section (Theorem 6.1) is still valid if we require further that all $\lambda_{i},i=1,\ldots,m,$ are nonnegative.
(ii) Non-degenerate mappings have a number of remarkable properties which make them an attractive domain for various applications; see [15, 17, 24, 25, 38]. Furthermore, the class of polynomial mappings (with fixed Newton polyhedra), which are non-degenerate at infinity, is generic in the sense that it is an open and dense semi-algebraic set; see [14, Theorem 1.2].
6.2. Necessary and sufficient conditions for exact penalization
In this subsection, let $f,g_{1},\ldots,g_{m}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ be polynomial functions and
$S$ $:=$ $\{x\in\mathbb{R}^{n}\mid g_{1}(x)\leq 0,\ldots,g_{m}(x)\leq 0\}.$
Let $\psi\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ be the semi-algebraic function defined by
$\psi(x)$ $:=$ $\max\{g_{1}(x),\ldots,g_{m}(x),0\},$
which is a residual function for $S.$ The first result of this section reads as follows.
Theorem 6.1.
Assume that the mapping $(f,g_{1},\ldots,g_{m})\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m+1}$ is convenient and non-degenerate at infinity. Then the following properties are equivalent:
  1. (i)
    there exist constants $c_{*}>0,\alpha>0,$ and $\beta>0$ with $\alpha\leq 1\leq\beta$ such that for all $c>c_{*},$
    $\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\left\{f(x)+c\left([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\right)\right\}\ >\ -\infty;$
  2. (ii)
    for any $M\geq 0,$ the polynomial $f$ is bounded from below on the set
    $S_{M}$ $:=$ $\{x\in\mathbb{R}^{n}\mid g_{1}(x)\leq M,\ldots,g_{m}(x)\leq M\}.$
When these equivalent properties hold, one has moreover for all $c>c_{*}$ that
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\left\{f(x)+c\left([\psi(x)]^{\alpha}+[\psi(x)]^{\beta}\right)\right\}\ \neq\ \emptyset.$
Proof.
(i) $\Rightarrow$ (ii) The implication is trivial.
(ii) $\Rightarrow$ (i) According to Theorem 5.3, it suffices to show that for all $M\geq 0,$ the restriction of $f$ on $S_{M}$ is coercive.
By contradiction, there is a real number $M\geq 0$ such that the restriction of $f$ on $S_{M}$ is not coercive. Certainly, the set $S_{M}$ is unbounded semi-algebraic, and so $S_{M}\cap\mathbb{S}_{r}$ is a nonempty compact set for all $r>R$ with $R$ being some positive constant (see [39, Corollary 2.11]). By Theorem 2.1, it is not hard to see that the function
$(R,\infty)\to\mathbb{R},\quad r\mapsto\min_{x\in S_{M},\|x\|=r}f(x),$
is semi-algebraic, and so is either constant or strictly monotone when $R$ is sufficiently large (in view of Lemma 2.12). Therefore, the limit
$\lim_{r\to\infty}\min_{x\in S_{M},\|x\|=r}f(x)$
exists and is finite, say, $\mathfrak{m}.$ On the other hand, the Fritz-John optimality conditions (see, for example, [2]) imply that the (semi algebraic) set
$\mathscr{A}\ :=\ \big\{(x,\lambda,\mu)\in\mathbb{R}^{n}\times\mathbb{R}^{m+1}\times\mathbb{R}$ $\mid$ $g_{i}(x)\leq M\textrm{ and }\lambda_{i}(g_{i}(x)-M)=0\textrm{ for }i=1,\ldots,m,$
$\ \lambda_{i}\geq 0\textrm{ for }i=0,1,\ldots,m,\ \|(\lambda,\mu)\|=1,\textrm{and}$
$\sum_{i=0}^{m}\lambda_{i}\nabla g_{i}(x)+\mu x=0\big\}$
is nonempty and unbounded. Here and in the following, it is convenient to write $\lambda:=(\lambda_{0},\lambda_{1},\ldots,\lambda_{m})$ and $g_{0}:=f.$ By applying Lemma 2.10 to the semi-algebraic function $\mathscr{A}\rightarrow\mathbb{R},(x,\lambda,\mu)\mapsto g_{0}(x),$ we get a smooth semi-algebraic curve
$(\varphi,\lambda,\mu)\colon(0,\epsilon)\rightarrow\mathbb{R}^{n}\times\mathbb{R}^{m+1}\times\mathbb{R},\quad t\mapsto(\varphi(t),\lambda(t),\mu(t)),$
satisfying the following conditions
  1. (c1)
    $\lim_{t\to 0^{+}}\|\varphi(t)\|=\infty;$
  2. (c2)
    $\lim_{t\to 0^{+}}g_{0}(\varphi(t))=\mathfrak{m};$
  3. (c3)
    $g_{i}(\varphi(t))\leq M$ for $i=1,\ldots,m;$
  4. (c4)
    $\lambda_{i}(t)\big(g_{i}(\varphi(t))-M\big)\equiv 0,$ for $i=1,\ldots,m;$
  5. (c5)
    $\lambda_{i}(t)\geq 0$ for $i=0,1,\ldots,m;$
  6. (c6)
    $\|(\lambda(t),\mu(t))\|\equiv 1;$
  7. (c7)
    $\sum_{i=0}^{m}\lambda_{i}(t)\nabla g_{i}(\varphi(t))+\mu(t)\varphi(t)\equiv 0.$
Since the (smooth) functions $\lambda_{i}$ and $g_{i}\circ\varphi$ are semi-algebraic, by shrinking $\epsilon$ if necessary, we can assume that these functions are either constant or strictly monotone (see Lemma 2.12). Then, by the condition (c4), one can see for all $i=1,\ldots,m$ that either $\lambda_{i}(t)\equiv 0$ or $g_{i}\circ\varphi(t)\equiv M,$ and hence
$\lambda_{i}(t)\frac{d}{dt}(g_{i}\circ\varphi)(t)$ $\equiv$ $0\quad\textrm{ for }\quad i=1,\ldots,m.$
It follows from the condition (c7) that
$\frac{\mu(t)}{2}\frac{d\|\varphi(t)\|^{2}}{dt}$ $=$ $\mu(t)\left\langle\varphi(t),\frac{d\varphi(t)}{dt}\right\rangle$
$=$ $-\sum_{i=0}^{m}\lambda_{i}(t)\left\langle\nabla g_{i}(\varphi(t)),\frac{d\varphi(t)}{dt}\right\rangle$
$=$ $-\sum_{i=0}^{m}\lambda_{i}(t)\frac{d}{dt}(g_{i}\circ\varphi)(t)$
$=$ $-\lambda_{0}(t)\frac{d}{dt}(g_{0}\circ\varphi)(t).$
This, together with the condition (c1), implies that $\lambda(t)\not\equiv 0$ because otherwise $\mu(t)\equiv\lambda_{0}(t)\equiv 0,$ which contradicts the condition (c6).
Let $J:=\{j\in\{1,\ldots,n\}\mid\varphi_{j}(t)\not\equiv 0\}\neq\emptyset.$ By Lemma 2.11, for each $j\in J,$ we can expand the coordinate $\varphi_{j}$ as follows
$\varphi_{j}(t)={x}^{0}_{j}t^{q_{j}}+\cdots,$
where ${x}^{0}_{j}\neq 0$, $q_{j}\in\mathbb{Q},$ and the dots stand for the higher order terms in $t.$ Observe from the condition (c1) that $\min_{j\in J}q_{j}<0.$
Let $q_{j}:=M^{\prime}$ for $j\not\in J$ with $M^{\prime}$ being sufficiently large and satisfying
$M^{\prime}$ $>$ $\max_{i=0,1,\ldots,m}\ \max_{\kappa\in\Gamma(g_{i})}\ \sum_{j\in J}q_{j}\kappa_{j}$
Take any $i\in\{0,1,\ldots,m\}$ and let $d_{i}$ be the minimal value of the linear function $\sum_{j=1}^{n}q_{j}\kappa_{j}$ on $\Gamma(g_{i})$ and let $\Delta_{i}$ be the maximal face of $\Gamma(g_{i})$ (maximal with respect to the inclusion of faces) where the linear function takes this value, i.e.,
$d_{i}$ $:=$ $d(q,\Gamma(g_{i}))\quad\textrm{ and }\quad\Delta_{i}\ :=\ \Delta(q,\Gamma(g_{i})).$
Since $g_{i}$ is convenient, $d_{i}<0$ and the restriction of $g_{i}$ on $\mathbb{R}^{J}$ is not constant; in particular, $\Gamma(g_{i})\cap{\mathbb{R}^{J}}=\Gamma(g_{i}|_{\mathbb{R}^{J}})$ is nonempty and different from $\{0\}.$ (Recall that $\mathbb{R}^{J}:=\{\kappa:=(\kappa_{1},\ldots,\kappa_{n})\in\mathbb{R}^{n}\mid\kappa_{j}=0\textrm{ for }j\not\in J\}.$) Furthermore, by definition of the vector $q,$ one has
$d_{i}$ $=$ $d(q,\Gamma(g_{i}|_{\mathbb{R}^{J}}))\quad\textrm{ and }\quad\Delta_{i}\ =\ \Delta(q,\Gamma(g_{i}|_{\mathbb{R}^{J}}))\ \subset\ {\mathbb{R}^{J}}.$
In particular, for each $j\not\in J,$ the polynomial $g_{i,\Delta_{i}}$ does not depend on the variable $x_{j}.$ Now suppose that $g_{i}$ is written as $g_{i}(x)=\sum_{\kappa}a_{i,\kappa}x^{\kappa}.$ We have
$g_{i}(\varphi(t))$ $=$ $\sum_{\kappa\in\Gamma(g_{i})\cap\mathbb{R}^{J}}a_{i,\kappa}(\varphi(t))^{\kappa}$
$=$ $\sum_{\kappa\in\Gamma(g_{i}|_{\mathbb{R}^{J}})}\Big(a_{i,\kappa}\prod_{j\in J}\varphi_{j}(t)^{\kappa_{j}}\Big)$
$=$ $\sum_{\kappa\in\Gamma(g_{i}|_{\mathbb{R}^{J}})}\Big(a_{i,\kappa}\prod_{j\in J}({x}^{0}_{j}t^{q_{j}})^{\kappa_{j}}+\cdots\Big)$
$=$ $\sum_{\kappa\in\Gamma(g_{i}|_{\mathbb{R}^{J}})}\Big(a_{i,\kappa}({x}^{0})^{\kappa}t^{\sum_{j\in J}q_{j}\kappa_{j}}+\cdots\Big)$
$=$ $\sum_{\kappa\in\Delta_{i}}a_{i,\kappa}({x}^{0})^{\kappa}t^{d_{i}}+\cdots$
where ${x}^{0}:=({x}^{0}_{1},\ldots,{x}^{0}_{n})\in(\mathbb{R}^{*})^{n}$ with ${x}^{0}_{j}:=1$ for $j\not\in J,$ and as usual, the dots stand for the higher order terms in $t.$ Recall that $g_{i,\Delta_{i}}({x})=\sum_{\kappa\in\Delta_{i}}a_{i,\kappa}x^{\kappa}.$ Hence
$g_{i}(\varphi(t))$ $=$ $g_{i,\Delta_{i}}({x}^{0})t^{d_{i}}+\cdots.$
On the the hand, we know that $d_{i}<0.$ Therefore, it follows from the conditions (c2) and (c4) that $g_{i,\Delta_{i}}({x}^{0})=0$ for $i=0$ or $i>0$ with $\lambda_{i}(t)\not\equiv 0.$ Then, by setting
$I$ $:=$ $\{i\in\{0,1,\ldots,m\}\mid\lambda_{i}(t)\not\equiv 0\},$
we get
$g_{i,\Delta_{i}}({x}^{0})$ $=$ $0\quad\textrm{ for all }\quad i\in I.$ (7)
(Note that $I\neq\emptyset$ because $\lambda(t)\not\equiv 0.$)
For $i\in I,$ expand the coordinate $\lambda_{i}$ in terms of the parameter (cf. Lemma 2.11) as follows
$\lambda_{i}(t)=\lambda_{i}^{0}t^{\theta_{i}}+\cdots,$
where $\lambda_{i}^{0}\neq 0$ and $\theta_{i}\in\mathbb{Q}.$ By the condition (c5), then $\lambda_{i}^{0}>0,$ which explains Remark 6.2(i).
For $i\in I$ and $j\in J$, by some similar calculations as with $g_{i}(\varphi(t))$, we get
$\frac{\partial g_{i}}{\partial x_{j}}(\varphi(t))$ $=$ $\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})t^{d_{i}-q_{j}}+\cdots.$
Consequently, we obtain for all $j\in J,$
$\sum_{i\in I}\lambda_{i}(t)\frac{\partial g_{i}}{\partial x_{j}}(\varphi(t))$ $=$ $\sum_{i\in I}\left(\lambda_{i}^{0}\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})t^{d_{i}+\theta_{i}-q_{j}}+\cdots\right)$ (8)
$=$ $\left(\sum_{i\in I^{\prime}}\lambda_{i}^{0}\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})\right)t^{\ell-q_{j}}+\cdots,$
where $\ell:=\min_{i\in I}(d_{i}+\theta_{i})$ and $I^{\prime}:=\{i\in I\mid d_{i}+\theta_{i}=\ell\}\neq\emptyset.$
Assume that we have proved:
$\sum_{i\in I^{\prime}}\lambda_{i}^{0}\nabla g_{i,\Delta_{i}}({x}^{0})$ $=$ $0.$ (9)
This equality, of course, when combined with (7) implies that the mapping $(g_{i})_{i\in I^{\prime}}\colon\mathbb{R}^{n}\to\mathbb{R}^{\#I^{\prime}}$ is not Khovanskii non-degenerate at infinity, a contradiction.
So we are left with proving (9). Indeed, for each $j\not\in J,$ the polynomial $g_{i,\Delta_{i}}$ does not depend on the variable $x_{j}$, so $\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}\equiv 0.$ Consequently,
$\sum_{i\in I^{\prime}}\lambda_{i}^{0}\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})$ $=$ $0\quad\textrm{ for all }\quad j\not\in J.$
If $\mu(t)\equiv 0$ then the condition (c7) and (8) give
$\sum_{i\in I^{\prime}}\lambda_{i}^{0}\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})$ $=$ $0\quad\textrm{ for all }\quad j\in J,$
and there is nothing to prove. So assume that $\mu(t)\not\equiv 0.$ By Lemma 2.11, we may write
$\mu(t)$ $=$ $\mu^{0}t^{\delta}+\cdots,$
where $\mu^{0}\neq 0$ and $\delta\in\mathbb{Q}.$ We deduce from the condition (c7) and (8) that $\ell-q_{j}\leq\delta+q_{j}$ for all $j\in J.$ Let $J^{\prime}:=\{j\in J\mid\ell-q_{j}=\delta+q_{j}\}.$ Assume that $J^{\prime}\neq\emptyset.$ Then
$\sum_{i\in I^{\prime}}\lambda_{i}^{0}\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})$ $=$ $\begin{cases}-\mu^{0}{x}^{0}_{j}&\textrm{ if }j\in J^{\prime},\\ 0&\textrm{ otherwise.}\end{cases}$
Hence
$\sum_{j=1}^{n}\left(\sum_{i\in I^{\prime}}\lambda_{i}^{0}\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})\right)q_{j}{x}^{0}_{j}$ $=$ $\sum_{j\in J^{\prime}}\left(\sum_{i\in I^{\prime}}\lambda_{i}^{0}\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})\right)q_{j}{x}^{0}_{j}$
$=$ $\sum_{j\in J^{\prime}}-q_{j}\mu^{0}({x}^{0}_{j})^{2}$
$=$ $-\frac{\ell-\delta}{2}\mu^{0}\sum_{j\in J^{\prime}}({x}^{0}_{j})^{2}.$
On the other hand, $g_{i,\Delta_{i}}$ is a weighted homogeneous polynomial of type $(q,d_{i}).$ Thus, from (6) and (7) we obtain for all $i\in I^{\prime},$
$\sum_{j=1}^{n}q_{j}{x}^{0}_{j}\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})$ $=$ $d_{i}\cdot g_{{}_{i},\Delta_{i}}({x}^{0})\ =\ 0.$
But $\ell-\delta\neq 0$ because $\ell-q_{j}\leq\delta+q_{j}$ for all $j\in J$ and $\min_{j\in J}q_{j}<0.$ Hence
$0$ $=$ $\sum_{i\in I^{\prime}}\left(\sum_{j=1}^{n}q_{j}{x}^{0}_{j}\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})\right)\lambda^{0}_{i}\ =\ \sum_{j=1}^{n}\left(\sum_{i\in I^{\prime}}\lambda^{0}_{i}\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})\right)q_{j}{x}^{0}_{j}$
$=$ $-\frac{\ell-\delta}{2}\mu^{0}\sum_{j\in J^{\prime}}({x}^{0}_{j})^{2}\ \neq\ 0,$
which is impossible. Therefore, $J^{\prime}=\emptyset,$ and so (9) holds. The theorem is proved. ∎
The second result of this section says that the exponent $\beta$ in Theorem 6.1 can be taken to be $1.$
Theorem 6.2.
Assume that the mapping $(f,g_{1},\ldots,g_{m})\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m+1}$ is convenient and non-degenerate at infinity. Then the following properties are equivalent:
  1. (i)
    there exist constants $c_{*}>0$ and $\alpha\in(0,1]$ such that for all $c>c_{*},$
    $\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\left\{f(x)+c\left([\psi(x)]^{\alpha}+\psi(x)\right)\right\}\ >\ -\infty;$
  2. (ii)
    there exists a constant $c_{*}>0$ such that
    $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c_{*}\psi(x)\big\}$ $>$ $-\infty.$
When these equivalent properties hold, one has moreover for all $c>c_{*}$ that
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\left\{f(x)+c\left([\psi(x)]^{\alpha}+\psi(x)\right)\right\}\ \neq\ \emptyset.$
Proof.
We note from [15, Theorem 1.1] that the set $\mathrm{argmin}_{S}f$ is nonempty when $f$ is bounded from below on $S.$
(i) $\Rightarrow$ (ii) The implication is trivial.
(ii) $\Rightarrow$ (i) We first show that there exist constants $c_{1}>0$ and $R>0$ such that
$c_{1}\psi(x)$ $\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad\|x\|\geq R.$ (10)
Indeed, if this were not true, there would exist a sequence $x_{k}\in\mathbb{R}^{n}$ with $\|x_{k}\|\to\infty$ such that
$k\psi(x_{k})$ $<$ $[f_{*}-f(x_{k})]_{+}\quad\textrm{ for all }\quad k.$
Then by a similar argument to the one given in the proof of Theorem 4.1, we can find a real number $\mathfrak{m}\in\mathbb{R}$ and sequences $x_{k}^{\prime}\in\mathbb{R}^{n}$ and $v_{k}\in\partial\psi(x_{k}^{\prime})$ such that $\|x_{k}^{\prime}\|\to\infty,$ $f(x_{k}^{\prime})\to\mathfrak{m},$ $\psi(x_{k}^{\prime})>0,$ $\psi(x_{k}^{\prime})\to 0,$ and
$\|x_{k}^{\prime}\|\left\|\frac{1}{{1+k}}\nabla f(x_{k}^{\prime})+\frac{k}{{1+k}}v_{k}\right\|$ $\to$ $0.$
For simplicity of notation, let $g_{0}:=f.$ According to Lemmas 2.4 and 2.10, there is a smooth semi-algebraic curve
$(\varphi,\lambda)\colon(0,\epsilon)\rightarrow\mathbb{R}^{n}\times\mathbb{R}^{m+1},\quad t\mapsto(\varphi(t),\lambda(t)),$
satisfying the following conditions
  1. (c1)
    $\lim_{t\to 0^{+}}\|\varphi(t)\|=\infty;$
  2. (c2)
    $\lim_{t\to 0^{+}}g_{0}(\varphi(t))=\mathfrak{m};$
  3. (c3)
    $\psi(\varphi(t))>0$ and $\lim_{t\to 0^{+}}\psi(\varphi(t))=0;$
  4. (c4)
    $\lambda_{i}(t)\big(g_{i}(\varphi(t))-\psi(\varphi(t))\big)\equiv 0$ for $i=1,\ldots,m;$
  5. (c5)
    $\lambda_{i}(t)\geq 0$ for $i=0,1,\ldots,m;$
  6. (c6)
    $\|\lambda(t)\|\equiv 1;$
  7. (c7)
    $\lim_{t\to 0^{+}}\|\varphi(t)\|\|\sum_{i=0}^{m}\lambda_{i}(t)\nabla g_{i}(\varphi(t))\|=0.$
Let $J:=\{j\in\{1,\ldots,n\}\mid\varphi_{j}(t)\not\equiv 0\}\neq\emptyset.$ By Lemma 2.11, for each $j\in J,$ we can expand the coordinate $\varphi_{j}$ as follows
$\varphi_{j}(t)={x}^{0}_{j}t^{q_{j}}+\cdots,$
where ${x}^{0}_{j}\neq 0$, $q_{j}\in\mathbb{Q},$ and the dots stand for the higher order terms in $t.$ Observe from the condition (c1) that $\min_{j\in J}q_{j}<0.$
Next, define $q,d_{i}$ and $\Delta_{i}$ for $i=0,1,\ldots,m,$ as in the proof of Theorem 6.1. Since $g_{i}$ is convenient, $d_{i}<0$ and the restriction of $g_{i}$ on $\mathbb{R}^{J}$ is not constant. Furthermore, by definition of the vector $q,$ one has
$d_{i}$ $=$ $d(q,\Gamma(g_{i}|_{\mathbb{R}^{J}}))\quad\textrm{ and }\quad\Delta_{i}\ =\ \Delta(q,\Gamma(g_{i}|_{\mathbb{R}^{J}}))\ \subset\ {\mathbb{R}^{J}}.$
In particular, for each $j\not\in J,$ the polynomial $g_{i,\Delta_{i}}$ does not depend on the variable $x_{j}.$ A direct calculation shows that
$g_{i}(\varphi(t))$ $=$ $g_{i,\Delta_{i}}({x}^{0})t^{d_{i}}+\cdots.$
where ${x}^{0}:=({x}^{0}_{1},\ldots,{x}^{0}_{n})\in(\mathbb{R}^{*})^{n}$ with ${x}^{0}_{j}:=1$ for $j\not\in J,$ and as usual, the dots stand for the higher order terms in $t.$ It follows from the conditions (c2), (c3), and (c4) that $g_{i,\Delta_{i}}({x}^{0})=0$ for $i=0$ or $i>0$ with $\lambda_{i}(t)\not\equiv 0.$ Then, by setting
$I$ $:=$ $\{i\in\{0,1,\ldots,m\}\mid\lambda_{i}(t)\not\equiv 0\}\ \neq\ \emptyset,$
we get
$g_{i,\Delta_{i}}({x}^{0})$ $=$ $0\quad\textrm{ for all }\quad i\in I.$
For $i\in I,$ expand the coordinate $\lambda_{i}$ in terms of the parameter (cf. Lemma 2.11) as follows
$\lambda_{i}(t)=\lambda_{i}^{0}t^{\theta_{i}}+\cdots,$
where $\lambda_{i}^{0}>0$ and $\theta_{i}\in\mathbb{Q}.$ From the condition (c6) one has $\theta_{i}\geq 0$ for all $i\in I$ with the equality occurring for some $i\in I.$
For $i\in I$ and $j\in J$, we have
$\frac{\partial g_{i}}{\partial x_{j}}(\varphi(t))$ $=$ $\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})t^{d_{i}-q_{j}}+\cdots.$
It implies for each $j\in J$ that
$\sum_{i\in I}\lambda_{i}(t)\frac{\partial g_{i}}{\partial x_{j}}(\varphi(t))$ $=$ $\left(\sum_{i\in I^{\prime}}\lambda_{i}^{0}\frac{\partial g_{i,\Delta_{i}}}{\partial x_{j}}({x}^{0})\right)t^{\ell-q_{j}}+\cdots,$
where $\ell:=\min_{i\in I}(d_{i}+\theta_{i})$ and $I^{\prime}:=\{i\in I\mid d_{i}+\theta_{i}=\ell\}\neq\emptyset.$ Note that $\ell<0$ because we know that $d_{i}<0$ for all $i=0,1,\ldots,m,$ and $\theta_{i}=0$ for some $i\in I.$ Thus, it follows from the condition (c7) that
$\sum_{i\in I^{\prime}}\lambda_{i}^{0}\nabla g_{i,\Delta_{i}}({x}^{0})$ $=$ $0.$
By definition, then the mapping $(g_{i})_{i\in I^{\prime}}\colon\mathbb{R}^{n}\to\mathbb{R}^{\#I^{\prime}}$ is not Khovanskii non-degenerate at infinity, which contradicts our assumption. Therefore, (10) holds.
Next, by applying Corollary 5.1 to the semi-algebraic functions $\phi(x):=[f_{*}-f(x)]_{+}$ and $\psi$ on the ball $\mathbb{B}_{R},$ we get constants $c_{2}>0$ and $\alpha\in(0,1]$ such that
$c_{2}[\psi(x)]^{\alpha}$ $\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad\|x\|\leq R.$
Letting $c_{*}:=\max\{c_{1},c_{2}\},$ we obtain
$c_{*}\big([\psi(x)]^{\alpha}+\psi(x)\big)$ $\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad x\in\mathbb{R}^{n}.$
This, together with Theorem 3.1, proves the implication (ii) $\Rightarrow$ (i). ∎
It is well known (see [8, Corollary 2.4.1] and [26, Theorem 4.4]) that if the (MFCQ) is satisfied at a local minimizer $\overline{x}$ to (P), then there exists a constant $c_{*}>0$ such that for all $c>c_{*},$ $\overline{x}$ is a local minimizer to the unconstrained optimization problem
$\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c\psi(x)\big\}.$
The following theorem is inspired by this result.
Theorem 6.3.
Let the mapping $(f,g_{1},\ldots,g_{m})\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m+1}$ be convenient and non-degenerate at infinity. If the (MFCQ) holds at every global minimum of $f$ on $S,$ then the following properties are equivalent:
  1. (i)
    there exists a constant $c_{*}>0$ such that for all $c>c_{*},$
    $\inf_{x\in S}f(x)$ $=$ $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c\psi(x)\big\}\ >\ -\infty;$
  2. (ii)
    there exists a constant $c_{*}>0$ such that
    $\inf_{x\in\mathbb{R}^{n}}\big\{f(x)+c_{*}\psi(x)\big\}$ $>$ $-\infty.$
When these equivalent properties hold, one has moreover for all $c>c_{*}$ that
$\mathrm{argmin}_{S}f(x)$ $=$ $\mathrm{argmin}_{\mathbb{R}^{n}}\big\{f(x)+c\psi(x)\big\}\ \neq\ \emptyset.$
Proof.
(i) $\Rightarrow$ (ii) The implication is trivial.
(ii) $\Rightarrow$ (i) The proof of Theorem 6.2 ensures the existence of constants $c_{1}>0$ and $R>0$ such that
$c_{1}\psi(x)$ $\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad\|x\|\geq R.$
Now, by a similar argument to the one given in the proof of Theorem 4.1, we can find a constant $c_{2}>0$ such that
$c_{2}\psi(x)$ $\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad\|x\|\leq R.$
Letting $c_{*}:=\max\{c_{1},c_{2}\},$ we obtain
$c_{*}\psi(x)$ $\geq$ $[f_{*}-f(x)]_{+}\quad\textrm{ for all }\quad x\in\mathbb{R}^{n}.$
Finally, the implication (ii) $\Rightarrow$ (i) follows directly from Theorem 3.1. ∎
Remark 6.3.
For simplicity of presentation we did not take equality constraints into account, but the results could be extended to both equality and inequality constraints.
7. Conclusion
In this paper, we have presented necessary and sufficient conditions for the exactness of penalty functions in optimization problems whose constraint sets are not necessarily bounded. The conditions are given in terms of properties of the objective and residual functions of the problems in question. Our results extend and improve some known facts in the literature on exact penalty functions. It would be nice to have applications of the results given here in optimization. This will be studied in the future research.
Acknowledgments
A part of this work was done while the second author and the third author were visiting Academy for Advanced Interdisciplinary Studies, Northeast Normal University, Changchun, China. The last version of the paper was completed when the first author and the second author were visiting the Vietnam Institute for Advanced Study in Mathematics (VIASM). The authors would like to thank these organizations for their hospitality and support.
Data availability
There is no data included in this paper.
Declarations
No potential conflict of interest was reported by the authors.